### v1

# On Subjective Uncertainty Quantification and Calibration in Natural Language Generation

Abstract
Applications of large language models often involve the generation of free-form responses, in which case uncertainty quantification becomes challenging. This is due to the need to identify task-specific uncertainties (e.g., about the semantics) which appears difficult to define in general cases. This work addresses these challenges from a perspective of Bayesian decision theory, starting from the assumption that our utility is characterized by a similarity measure that compares a generated response with a hypothetical true response. We discuss how this assumption enables principled quantification of the model’s subjective uncertainty and its calibration. We further derive a measure for epistemic uncertainty, based on a missing data perspective and its characterization as an excess risk. The proposed measures can be applied to black-box language models. We demonstrate the proposed methods on question answering and machine translation tasks, where they extract broadly meaningful uncertainty estimates from GPT and Gemini models and quantify their calibration.
1 Introduction
We are interested in uncertainty quantification for language models (LMs) in free-form natural language generation (NLG): given instruction $I$, the model generates a response $y^{\prime}$ based on its predictive distribution $p_{M}(y\mid I)$, where $y^{\prime}$ could be any natural language passage that fits the instruction. An example of this is question answering (QA): given a question from the user, the model may provide a brief answer, but it may also follow with supporting facts and explanations, which can vary in form and detail. The user can be satisfied by a wide variety of responses, irrespective of their style or (to some extent) the choice of supporting facts included.
Free-form NLG poses significant challenges to uncertainty quantification: some aspects of generation are irrelevant to the task’s purpose and best excluded from uncertainty quantification, but it often appears that we are unable to characterize them precisely. If left unaddressed, however, the model’s variation in the irrelevant aspects may dominate in standard uncertainty measures such as token-level entropy (Kuhn et al.,, 2023), making them uninformative about the model’s actual performance on the task.
Starting from Kuhn et al., (2023), a recent line of work (Kuhn et al.,, 2023; Lin et al.,, 2024; Zhang et al.,, 2023; Aichberger et al.,, 2024) studied this issue and proposed measuring the “semantic uncertainty” of generation; “semantics” is defined as the equivalence class of textual responses that logically entail one another. Empirical improvements in downstream tasks evidenced their contributions and highlighted the importance of task-specific uncertainty quantification, but important conceptual and practical issues remain. From a practical perspective, semantic equivalence is estimated using machine learning models, resulting in imprecise estimates that do not necessarily define an equivalence relation. The imprecision necessitates the introduction of heuristics to post-process the estimates or to aggregate them through other means. Conceptually, the notion of semantic equivalence does not always provide a valid or complete characterization of relevance for certain tasks. Style transfer (Jin et al.,, 2022) tasks provide an example of the first kind. For the second scenario, consider a QA task where the model is fully certain about the answer to a question, as well as a large set of supporting facts, each of which provides complete and independent justification for the answer; yet there is still the “uncertainty”, or variation across generations, about which facts are included in a response. Then very few of the model generations may logically entail one another, and there is a very high level of “semantic uncertainty” just as if there was complete uncertainty about the correct answer.
This paper is about the observation that the above challenges can be resolved in a more general setup from a perspective of Bayesian decision theory (Savage,, 1954). We first assume that we can compute a similarity-type measure $S(y^{\prime},y;I)$ that characterizes our utility function when the model generates $y^{\prime}$ in respond to $I$ and $y$ is a (true or hypothetical) correct response. This generalizes previous works based on semantic equivalence, which corresponds to defining $S$ using entailment. It can also be applied to traditional structured prediction tasks where a choice of $S$ is readily available (e.g., the chrF score for machine translation, Popović,, 2015), as well as tasks where evaluation can be adequately implemented using frontier language models that computes $S$ based on few-shot demonstrations or detailed instructions.
To quantify the subjective uncertainty defined by the LM, we assume the generation $y$ is always chosen to maximize the expected utility under the LM’s predictive distribution $p_{M}$2.1). This is a common assumption in NLG (Bickel and Doksum,, 1997; Bertsch et al.,, 2023). It can be supported by a common belief that high-capacity LMs may approximate Bayesian inference (e.g., Xie et al.,, 2021; Akyürek et al.,, 2022; Hahn and Goyal,, 2023; Ye et al.,, 2024), and is also more broadly applicable whenever we view $p_{M}$ as our best available model for $y\mid I$. Subjective uncertainty is then naturally characterized by the Bayes risk, or equivalently the maximum achievable expected utility, where the action space is defined by the candidate generations available. This simple observation allows us to understand that previous methods for “semantic uncertainty” can be adapted to a broader range of scenarios, regardless of whether semantic equivalence is relevant. It also provides a unique, principled approach for aggregating similarity measures among generations (§2.2).
For subjective uncertainty measures to be useful, it will be ideal if the LM is calibrated. It has been unclear in previous works how calibration can be evaluated in free-form NLG tasks, given the distinction between relevant and irrelevant differences for a specific task and our apparent inability to formally characterise it. The decision-theoretic view provides a natural answer to this question: an LM is deemed as to possess a calibrated notion of uncertainty if the expected subjective utility of its actions matches the expectation of the actually incurred utility, defined w.r.t. responses from the true data distribution (§2.3). This observation allows us to quantify the calibration of LMs through reliability diagrams (Murphy and Winkler,, 1977) and a generalized version of expected calibration error (Naeini et al.,, 2015). In the discussion we also note a generally applicable definition for “task-specific semantics”: the equivalence class defined by the action-utility map, known as Savage acts.
Bayesian modelling allows for the decomposition of predictive uncertainty into epistemic uncertainty and aleatoric uncertainty (Der Kiureghian and Ditlevsen,, 2009). The decomposition allows us to better understand and diagnose the model’s uncertainty (e.g., only the epistemic uncertainty should decrease as the training set grows); it can also guide data acquisition (Kendall and Gal,, 2017). The quantification of epistemic uncertainty in general NLG applications appears challenging, due to the black-box nature of LMs and the free-form generations involved. §2.4 demonstrates that the decision-theoretic view enables us to quantify epistemic uncertainty for in-context learning (ICL, Brown et al.,, 2020) applications; key additional ingredients include a missing data perspective to Bayesian modelling (Fong et al.,, 2024) and a connection between epistemic uncertainty and a notion of excess risk (Xu and Raginsky,, 2022). The resulted measure is connected to and generalizes various previous works on epistemic uncertainty quantification for black-box models, which provides independent justification for its use.
The rest of this paper is structured as follows: §2 presents the proposed methodology and discusses connections to previous work; §3 provides experimental illustrations; and §4 provides concluding remarks.
2 Quantifying Subjective Uncertainty and Calibration
2.1 A Utilitarian Setup
The basic setup of this paper is as follows:
  • we have a NLG task: given prompt $I\in\mathcal{I}$, generate a response $y^{\prime}\in\mathcal{Y}$ where $\mathcal{Y}$ denotes the space of natural language responses;
  • our utility can be measured through the similarity measure $S(y^{\prime},y;I)\in\mathbb{R}$, which can be cheaply evaluated;
  • we have access to an LM $p_{M}$, and, in the absence of further evidence, consider $p_{M}$ sufficiently trustworthy so that the ideal generation (i.e., action) $y^{\prime}$ given $I$ should maximize the expected utility,
    $y^{\prime}:=\operatorname*{arg\,max}_{y^{\prime}\in\mathcal{Y}^{\prime}_{I}}\mathbb{E}_{y\sim p_{M}(\cdot\mid I)}S(y^{\prime},y;I),$ (1)
    where $\mathcal{Y}^{\prime}_{I}\subset\mathcal{Y}$ denotes the space of candidate generations.
In the above, $p_{M}$ represents a subjective belief about the true data distribution. $\mathcal{Y}_{I}^{\prime}\subset\mathcal{Y}$ can be determined based on computational constraints, e.g., as a separate set of samples from $p_{M}(\cdot\mid I)$.111 When $\mathcal{Y}_{I}^{\prime}$ is stochastic, expected utility/risk quantities such as (1) should also be averaged over the randomness in $\mathcal{Y}_{I}^{\prime}$. For brevity, we omit such (outmost) expectations throughout the paper. The utility $S$ determines a risk function $r(y^{\prime},y;I)=-S(y^{\prime},y;I)$, which can be scaled and shifted as desired.
As discussed in §1, this setup is very general. This decision-theoretic perspective is taken in the framework of minimum Bayes risk (MBR) decoding (Bickel and Doksum,, 1997), which advocates for generating $y$ based precisely on (1). MBR recovers many modern methods for natural language generation (Bertsch et al.,, 2023), and these methods can be viewed as having adopted the same view implicitly. Note that adopting the decision-theoretic setup does not compel us to assume that $p_{M}$ is a strictly Bayesian model (e.g., in the sense of Xie et al.,, 2021): although that will provide a good justification, it is also possible that a non-Bayesian LM may better approximate our belief than any available Bayesian model does.
2.2 Task-Specific Measure of Subjective Uncertainty
It is well known that Bayesian uncertainty measures are connected to minimum achievable risks of the form
$R_{B}(I;p_{M},\mathcal{Y}^{\prime}_{I}):=\min_{y^{\prime}\in\mathcal{Y}^{\prime}_{I}}\mathbb{E}_{y\sim p_{M}(\cdot\mid I)}r(y^{\prime},y;I).$ (2)
For example, if $\mathcal{Y}=\mathcal{Y}^{\prime}_{I}\subset\mathbb{R}$ and $r(y,y^{\prime};I)=(y-y^{\prime})^{2}$, $R_{B}(I;p_{M},\mathcal{Y}^{\prime}_{I})$ will be equivalent to the variance of the predictive distribution; if with an abuse of notation we redefine $y^{\prime}$ as a probability density function and $r(y^{\prime},y;I)\leftarrow\log y^{\prime}(y)$, $R_{B,\mathcal{Y}^{\prime}_{I}}(I)=H[p_{M}(\cdot\mid I)]$ will be the entropy. When the action space $\mathcal{Y}^{\prime}_{I}\subsetneq\mathcal{Y}$ is a strict subset, (2) will be connected to measures of usable information as discussed in Xu et al., (2019).
It is thus natural to also adopt (2) as a measure of subjective uncertainty for NLG tasks. To implement (2), we adopt Monte Carlo approximation for the expectation.
A common generation strategy is to follow the Gibbs predictor, which corresponds to (1) with $\mathcal{Y}^{\prime}_{I}:=\{y_{I}\}$ and $y_{I}\sim p_{M}(\cdot\mid I)$ being a single sample. In such cases, the expected uncertainty $\mathbb{E}_{\mathcal{Y}^{\prime}_{I}}R_{B}(I;p_{M},\mathcal{Y}^{\prime}_{I})$ is connected to the degree statistic in Lin et al., (2024, Eq. 8); the difference is that we use a similarity measure that is assumed to define our utility. While Lin et al., (2024) explored 9 different uncertainty measures for the Gibbs predictor, the decision-theoretic perspective leads to one unique method without any underspecified design choices, justifies its use beyond the scope of semantic equivalence, and highlights that different choices of predictors ($\mathcal{Y}_{I}^{\prime}$) should be matched with different uncertainty measures. However, being a measure of subjective uncertainty, it is not guaranteed to always improve on downstream tasks compared with heuristic alternatives, especially if the LM is less calibrated.
2.3 Evaluation of Task-Specific Calibration
Subjective uncertainty measures are most useful when the LM is calibrated. As discussed in Huang et al., (2024), many previous works studying uncertainty on NLG tasks did not distinguish between calibration and predictive performance in their evaluation, which leads to a relative lack of understanding about the calibration of LMs in such scenarios.
A typical definition of calibration is as follows (see e.g., Johnson et al.,, 2024): suppose the spaces $\mathcal{Y},\mathcal{I}$ are discrete, and denote the true data distribution as $p_{0}$; a model $p_{M}$ is considered calibrated w.r.t. some grouping function $G(I):\mathcal{I}\to\mathcal{G}$ if we have
$\mathbb{E}_{p_{0}}(p_{M}(y=y^{\prime}\mid I)\mid G(I)=g)=\mathbb{E}_{p_{0}}(p_{0}(y=y^{\prime}\mid I)\mid G(I)=g)~{}~{}\forall g\in\mathcal{G},y^{\prime}\in\mathcal{Y}.$ (3)
The grouping function $G$ may be defined based on the model’s prediction, or it may incorporate additional information such as input demographics. In the NLG setting, however, the criterion (3) can be trivially violated if, e.g., the LM always generates a longer response compared with $p_{0}$.
In light of the assumptions in §2.1, it is natural to relax (3) as follows:
$\mathbb{E}(p_{M}(\{y:S(y^{\prime},y;\mathcal{I})=s\}\mid I)\mid G(I)=g)$ $=\mathbb{E}(p_{0}(\{y:S(y^{\prime},y;\mathcal{I})=s\}\mid I)\mid G(I)=g)$
$\hskip 130.0002pt~{}~{}\forall g\in\mathcal{G},y^{\prime}\in\mathcal{Y}^{\prime}_{I},s\in\mathbb{R}.$ (3’)
(3) is equivalent to (3) if $\mathcal{Y}^{\prime}_{I}=\mathcal{Y}$ and $S(y^{\prime},y;\mathcal{I})=\mathbf{1}\{y=y^{\prime}\}$, or more generally if for all $I\in\mathcal{I}$ the map from $y$ to the function
$s_{y,I}:\mathcal{Y}_{I}^{\prime}\to\mathbb{R},~{}y^{\prime}\mapsto S(y^{\prime},y;I)$ (4)
is injective (e.g., if $S$ is defined by a metric on $\mathcal{Y}$). (3) becomes weaker if $S(y^{\prime},y;\mathcal{I})$ only depends on $y$ and $y^{\prime}$ through a non-invertible function, e.g., a map from a natural language passage to its semantics; in such cases, it is clear that (3) remains a sufficiently strong criterion for task-specific calibration. More generally, we can always view the function $s_{y,I}(y^{\prime})$ as a representation of “task-specific semantics”.222 For the goals of providing the best generation from $\mathcal{Y}^{\prime}_{I}$ and quantifying its risk, it suffices to define the function $s_{y,I}$ on $\mathcal{Y}^{\prime}_{I}$. For purposes beyond the scope of the generation task, it may be helpful to extend its domain to a larger subset of $\mathcal{Y}$. In fact, the replacement of $y^{\prime}\in\mathcal{Y}^{\prime}_{I}$ with the more compact $s_{y^{\prime},I}(\cdot)$ has been advocated by Savage, (1954), and $s_{y,I}$ is referred to as a Savage act (see e.g., Marinacci,, 2015, §2.4).
A main reason we are interested in criteria such as (3) or (3) is that they guarantee the calibration of uncertainty measures. It follows from (3) that $\mathbb{E}_{p_{0}(I\mid G(I)=g)p_{M}(y\mid I)}r(y^{\prime},y;I)=\mathbb{E}_{p_{0}(I\mid G(I)=g)p_{0}(y\mid I)}r(y^{\prime},y;I)$ for all $y^{\prime}$, and thus
$\mathbb{E}_{p_{0}(I\mid G(I)=g)}R_{B}(I;p_{M},\mathcal{Y}^{\prime}_{I})=\mathbb{E}_{p_{0}(I\mid G(I)=g)p_{M}(y\mid I)}r(\hat{y}_{I},y;I)=\mathbb{E}_{p_{0}(I\mid G(I)=g)p_{0}(y\mid I)}r(\hat{y}_{I},y;I),$ (5)
where $\hat{y}_{I}$ denotes the MBR generation (1). Plugging in $G(I)=R_{B}(I;p_{M},\mathcal{Y}^{\prime}_{I})$, we find that LMs are calibrated in the sense of (3) only if
$\text{gECE}(p_{M}):=\mathbb{E}_{s}|f_{M}(s)-s|=0,~{}~{}~{}\text{where }f_{M}(s):=\mathbb{E}_{I,y,\hat{y}_{I}}(r(\hat{y}_{I},y;I)\mid R_{B}(I;p_{M},\mathcal{Y}^{\prime}_{I})=s)~{}\forall s\in\mathbb{R}.$ (6)
(6) provides a natural generalization of the expected calibration error (ECE, Naeini et al.,, 2015) in classification, which it recovers with $r(y^{\prime},y;I)\leftarrow\mathbf{1}\{y^{\prime}\neq y\}$. Similar to the classification case, (6) can be estimated through histogram binning. The plot for the function $f_{M}$ is called a reliability diagram (Murphy and Winkler,, 1977) and provides information on overconfidence ($f_{M}(s)<s$) or underconfidence. Importantly, these methods remain applicable in NLG as we have shown. Moreover, they will automatically focus on the task-specific aspects of uncertainty, since the function $f_{M}$ only depends on model generations through the utility $S$.
Remark 2.1.
We discussed the evaluation of LM calibration based on the uncertainty measure (2). To our knowledge, only Huang et al., (2024) studied a similar problem in the general NLG setting, namely the rank-calibration of general uncertainty measures. As discussed in Huang et al., (2024, §4.2), rank-calibration is still a different property; in particular, (2) will be rank-calibrated whenever $f_{M}$ is monotone, even though it can be arbitrarily overconfident or underconfident and thus miscalibrated.
2.4 Representing and Eliciting Epistemic Uncertainty
We now turn to the problem of epistemic uncertainty quantification in in-context learning, where the prompt $I=\langle x_{1},y_{1},\ldots,x_{n},y_{n},x_{*}\rangle$ is formed by $n$ demonstrations $(x_{i},y_{i})$ followed by a test input $x_{*}$. We assume that $r(y,y^{\prime};I)=r(y,y^{\prime};x_{*})$ only depends on $I$ through $x_{*}$, which is commonly satisfied in applications. We propose the following measure of epistemic uncertainty inspired by Fong et al., (2024); Xu and Raginsky, (2022):
$\inf_{y^{\prime}\in\mathcal{Y}^{\prime}_{I}}\mathbb{E}_{p_{x}(x_{n+1:N})}\mathbb{E}_{p_{M}(y_{n+1:N}\mid y_{1:n},x_{1:N})}\mathbb{E}_{p_{M}(y_{*}\mid y_{1:N},x_{1:N},x_{*})}r(y^{\prime},y_{*};x_{*})$
$\hskip 40.00006pt-\mathbb{E}_{p_{x}(x_{n+1:N})}\mathbb{E}_{p_{M}(y_{n+1:N}\mid y_{1:n},x_{1:N})}\inf_{y^{\prime}\in\mathcal{Y}^{\prime}_{I}}\mathbb{E}_{p_{M}(y_{*}\mid y_{1:N},x_{1:N},x_{*})}r(y^{\prime},y_{*};x_{*}).$ (7)
In the above, $p_{x}$ denotes a user-specified distribution of additional ICL inputs, $y_{n+1:N}$ are model generations, $N>n$ is a hyperparameter, and we assume the LM has good performance for ICL up to a sample size of $N$. Eq. (7) can be estimated through autoregressive sampling. As long as we use the same set of samples for its both terms, the resulted estimate will always be non-negative, due to the convexity of the functional $\rho\mapsto\inf_{y^{\prime}\in\mathcal{Y}^{\prime}_{I}}\mathbb{E}_{\rho(y_{*})}r(y^{\prime},y_{*};x_{*})$.
To understand (7), first consider a simplified scenario as follows: suppose the LM $p_{M}$ is equivalent to a Bayesian model that assumes $\{(x_{i},y_{i})\}$ are i.i.d. conditional on a latent variable $\theta$, $p_{x}$ defines i.i.d. $x_{n+1:N}$, and $\{x_{n+i},y_{n+i}\}$ are sufficiently informative so that the likelihood function $p(y=\cdot\mid x=x_{*},\theta)$ determined by $\theta\mid x_{1:N},y_{1:N}$ is identifiable given infinite samples. Then under mild technical conditions, as $N\to\infty$ (7) will become equivalent to
$\min_{y^{\prime}\in\mathcal{Y}^{\prime}_{I}}\mathbb{E}_{p_{M}(\theta\mid x_{1:n},y_{1:n})}\mathbb{E}_{p(y\mid\theta,x=x_{*})}r(y^{\prime},y;x_{*})-\mathbb{E}_{p_{M}(\theta\mid x_{1:n},y_{1:n})}\min_{y^{\prime}\in\mathcal{Y}^{\prime}_{I}}\mathbb{E}_{p(y\mid\theta,x=x_{*})}r(y^{\prime},y;x_{*}).$ (7’)
The equivalence is a consequence of Doob, (1949)’s theorem and can be proved with the same idea as Fong et al., (2024). For completeness, we provide a proof in Appendix A. (7) is a subjective notion of excess risk. In other words, it is the amount of reducible risk should we have full knowledge about the distribution of $y\mid x=x_{*}$, which can be determined given either the latent $\theta$, or an infinite amount of missing data $\{x_{>n},y_{>n}\}$. As noted in Xu and Raginsky, (2022), this is precisely the intuition behind Bayesian epistemic uncertainty, and we recover standard epistemic uncertainty measures (e.g., posterior variance for a regression mean, or mutual information) if we return to simple scenarios where $r$ is the square or log loss.
Within the Bayesian framework, (7) with a finite choice of $N$ will provide a lower bound for (7): by convexity the second term may become larger. The intuition is that knowledge of the samples $\{x_{n+1:N},y_{n+1:N}\}$ does not always allow for the full reduction of risk. It is possible to view the lower bound as an indication of the “true epistemic uncertainty” (7), and it can be tightened if we choose $\{x_{n+i}\}$ to be more similar to $x_{*}$. The bound will also be tight regardless of the choice of $N$ if the uncertainty is “fully aleatoric” or “fully epistemic”, i.e., if the risk for $y\mid x=x_{*}$ cannot be reduced by any number of additional observations $(x_{n+1:N},y_{n+1:N})$ or can be fully reduced using a single $(x_{n+1},y_{n+1})$. Alternatively, we can observe that we do not always need the tightest bound for (7): if all we can do is to collect $N-n$ real samples, the remaining proportion of the risk will be effectively irreducible.
As discussed in §1, the Bayesian view above may be justified through a common belief that high-capacity LMs may approximate Bayesian inference. It is also supported by the results of Wen et al., (2022, §2.6, §4), which imply that in a multi-task setting, any $p_{M}$ with a strong average-case performance for prediction will yield a similar value for (7) as the true Bayesian model. Falck et al., (2024) investigated this belief empirically on synthetic low-dimensional datasets, and found it to be valid for smaller choices of $N$, although it breaks down as $N\to\infty$ becomes larger.333 Falck et al., (2024) also studied epistemic uncertainty quantification for ICL but restricted to datasets where the method of Fong et al., (2024) is applicable. Their method assumes $\{x_{n+1:N},$ $y_{n+1:N}\}$ can be modelled by a low-dimensional statistical model, so that a posterior for the model parameter can be constructed and visualized. This is unrealistic for free-form natural language data. Taken together, these results suggest that the Bayesian perspective could be relevant when we restrict $N$ to a scope where ICL is sample efficient.
Independent to the Bayesian perspective, Eq. (7) can also be understood through its connection to recent works. Collier et al., (2022) studied a broadly similar problem where a proportion of the combined uncertainty can be reduced by conditioning on privilege information; this leads to a similar measure for reducible uncertainty as (7), if we view $(x_{n+1:N},y_{n+1:N})$ as the latent “privilege information”. Johnson et al., (2024); Ahdritz et al., (2024) proposed methods that measure the correlation between consecutive predictions $(y_{n+1},y_{*})$ given the same input $x_{n+1}=x_{*}$; as argued in Johnson et al., (2024), any learning system $p_{M}$ with complete confidence for the true data distribution should lead to conditionally i.i.d. $(y_{*},y_{n+1})$ given $I$, and any $p_{M}$ with a “fully epistemic” uncertainty should produce $y_{*}=y_{n+1}$. It is clear that (7) matches these behaviours at the extremes: it equals 0 in the former case, and correctly indicate a full reduction of risk in the latter case (i.e., it equals $R_{B}(I;p_{M},\mathcal{Y}^{\prime}_{I})$ assuming $r(y,y;I)\equiv 0$).The Bayesian perspective provides additional insights to the works of Johnson et al., (2024); Ahdritz et al., (2024). Eq. (7) generalizes their methodology by allowing for a wider range of choices for $\{x_{n+1:N}\}$, and the above discussion provides guidance on their choices.
Finally, we note that while (7) is generally applicable to MBR decoding, with a trivial choice of $\mathcal{Y}_{I}^{\prime}$ it may take a trivial value: if we use the Gibbs predictor which always samples from $p_{M}(\cdot\mid x_{1:n},y_{1:n})$, we will find (7) to be zero; this is to be expected as the predictor does not make any attempt to utilize the additional information. In our experiments we choose $\mathcal{Y}_{I}^{\prime}$ to contain multiple samples, one from each sample path for $y_{n+1:N}$. As we will see, this is sufficient to provide meaningful estimates of epistemic uncertainty.
3 Experiments
3.1 Free-Form Question Answering
We first illustrate the proposed method on two QA tasks chosen by Lin et al., (2024): CoQA (Reddy et al.,, 2019) and NQOpen (Lee et al.,, 2019). CoQA contains reading comprehension questions where the answer can generally be deduced from the prompt, whereas NQOpen is an open-domain task and requires information not present in the prompt. Kalai and Vempala, (2024) proved in simplified scenarios that instruction-tuned LMs may have to be uncalibrated on the latter type of task; we investigate their theoretical predictions empirically.
Figure 1: Reliability diagrams for GPT-3.5 on QA tasks. Error bar denotes 95% confidence interval for the average observed utility.
We follow the setup in Lin et al., (2024) and adopt the GPT-3.5 model (OepnAI,, 2024), both for generation and to define the automated evaluation criterion; the latter determines our utility function and in turn the uncertainty measure (2).444 In this aspect our method differs from Lin et al., (2024), where uncertainty measures are defined with estimated entailment relations or lexical metrics. The evaluation criterion is defined using ICL, by instructing the LM to rate the consistency between two responses based on few-shot demonstrations (Lin et al.,, 2024, App. B.3). We reuse the model samples released by Lin et al., (2024), which contains 20 sampled responses for each query. Following their work we focus on the Gibbs predictor.
The results are shown in Fig. 1. As we can see, the LM’s subjective uncertainty generally correlates well with the average observed utility. It is slightly overconfident on the reading comprehension task, but much more so on the open-domain task. Such a difference appears consistent with the theoretical analysis of Kalai and Vempala, (2024).
3.2 In-Context Machine Translation
We now consider machine translation with ICL, where utility is defined through the widely adopted chrF score (Popović,, 2015). chrF measures both semantic and syntactic differences; the latter can be relevant in translation due to the need to detect unnatural or invalid syntactic structures. The inherent syntactic differences in the possible generations for $y\mid x$ constitute a source of irreducible uncertainty, making machine translation a non-trivial scenario to illustrate our uncertainty decomposition method.
We follow the setup of Agarwal et al., (2024) who demonstrated that ICL can lead to steady improvements for machine translation. We adopt the FLORES+ dataset (NLLB Team et al.,, 2022), in particular the translation tasks from English to the following languages: French, Tamil, Yue Chinese (Cantonese), and Tigrinya. We use the following LMs: GPT-3.5, Gemini-1.5-Flash, and Gemini-1.5-Pro (Gemini Team,, 2024), and vary the number of ICL shots $n\in\{4,16\}$. For each task, we subsample 128 queries $(x_{*},y_{*})$ from the datasets. For each query, we draw 35 samples from the LM to approximate its predictive distribution, and another 5 samples to define the action space $\mathcal{Y}^{\prime}_{I}$. For the quantification of epistemic uncertainty, we set $N=n+8$, and sample $x_{n+1:N}$ by instructing an LM to rewrite the test query $x_{*}$; see Appendix B for details.
$n$ gemini-1.5-flash gemini-1.5-pro gpt-3.5
$4$ 0.276 / 0.356 0.332 / 0.431 0.185 / 0.329
$16$ 0.246 / 0.371 0.304 / 0.426 0.168 / 0.351
Table 1: Machine translation (English $\to$ Yue Chinese): gECE (Eq. (6), $\downarrow$) and chrF ($\uparrow$) for varying choices of $n$.
We first evaluate the calibration of the combined subjective uncertainty on the Yue Chinese task. As shown in Table 1, the LMs are generally poorly calibrated, although calibration improves as the sample size $n$ increases. Calibration plots are presented in Figure 3 in Appendix, which shows that all LMs are overconfident. It is interesting to note an anticorrelation between the task performance and calibration among the three LMs evaluated. We note that Yue Chinese may constitute a low-resource translation task (Haddow et al.,, 2022) for which the LMs’ pretraining corpus is likely less informative. As is also shown in Figure 3, calibration is better on the French task for which more corpus is available.
Figure 2: Machine translation (English $\to$ Yue Chinese): reduction of epistemic uncertainty (EU) vs total uncertainty (TU) as the ICL sample size increases from $4$ to $16$.
We now investigate whether the LMs possess a coherent notion of epistemic and aleatoric uncertainty, in particular whether epistemic uncertainty is the only reducible component as sample size increases. For this purpose we plot the reduction of total uncertainty (TU, (2)) against that of epistemic uncertainty (EU, (7)) on each data point. Figure 2 shows that the two quantities have a good match. Deviation from perfect equivalence is expected here, due to the variance in the Monte-Carlo estimation for EU and TU.
Language French Tamil Yue Tigrinya
Avg. EU 0.016 0.059 0.078 0.091
chrF 0.737 0.392 0.329 0.084
# speakers (M) 312 87 87 10
Table 2: Machine translation: average epistemic uncertainty on different languages, obtained from gpt-3.5 using $n=4$ ICL shots. We also provide the chrF score and the number of speakers (from Ethnologue,, 2024, in million) of each language as an indication of its linguistic resource available.
Finally, we evaluate if the epistemic uncertainty measure (7) is meaningful, by comparing its average value across different translation tasks: we expect sensible notions of epistemic uncertainty to be higher on tasks for low-resource languages, as they are likely less represented in the pretraining corpus. Table 2 confirms this intuition: the LM demonstrates higher epistemic uncertainty on lower-resource languages.
4 Conclusion
This work studies uncertainty quantification in free-form natural language generation. Through a Bayesian decision-theoretic perspective, we derived principled methods to quantify the model’s task-specific subjective uncertainty and to evaluate its calibration. The proposed methods are also connected to previous work, which provides additional justification for their adoption. While the discussions (in particular those in §2.2 and §2.3) are not necessarily novel outside the context of language modelling, within this domain the decision-theoretic perspective appears new and addresses important conceptual challenges in uncertainty quantification. Experiments demonstrate that the methods extract meaningful uncertainty estimates from black-box LMs and quantify their calibration.
Our discussion concerns the quantification of subjective uncertainty. In applications where i.i.d. calibration samples are available, it would be beneficial to recalibrate (Marx et al.,, 2022) the uncertainty measures, which can be achieved through (6), or to incorporate the uncertainty measures in conformal prediction (Quach et al.,, 2023). As the capabilities of LMs continue to evolve, it is also conceivable that future-generation models may find use in “one-of-its-kind” applications where post-calibration and conformal approaches are inapplicable. In such scenarios, it would be crucial to have principled measures for subjective uncertainty, as well as the tools to dissect and understand the sources of this uncertainty, e.g., through the proposed decomposition into epistemic and aleatoric uncertainty. It would also be interesting for future work to investigate the consistency between the LM’s verbalized uncertainty (Kadavath et al.,, 2022) and the subjective uncertainty defined by its predictive distributions.
Appendix A Proof for the Equivalence between (7) and (7)
Claim A.1.
Suppose $\mathcal{X},\mathcal{Y}$ are finite sets, and $\sup_{y^{\prime},y,x}|r(y^{\prime},y;x)|<\infty$. Suppose there exists a Bayesian model with prior $\pi$ over a discrete parameter $\theta\in\Theta$ and a (conditional) likelihood function $p(y\mid x,\theta)$, s.t. for all $(n,x_{1:n},y_{1:n},x_{*})$ we have
$p_{M}(y_{n+1}=\cdot\mid x_{n+1}=x_{*},x_{1:n},y_{1:n})=\int\pi(d\theta\mid x_{1:n},y_{1:n})p(y=\cdot\mid x=x_{*},\theta)$
where $\pi(d\theta\mid x_{1:n},y_{1:n})\propto\pi(d\theta)\prod_{i=1}^{n}p(y=y_{i}\mid x=x_{i},\theta)$ denotes the parameter posterior. Suppose (7) is defined using $x_{n+1:N}\overset{i.i.d.}{\sim}p_{x,1}$, and that for all $x_{*}\in\mathcal{X}$ and $\pi$-a.e. $\theta$ we have
$\lim_{n\to\infty}\mathbb{E}_{x_{1:n}\sim p_{x,1},y_{i}\sim p(y=\cdot\mid x=x_{i},\theta)}\|p_{M}(y_{n+1}=\cdot\mid x_{n+1}=x_{*},x_{1:n},y_{1:n})-p(y=\cdot\mid x=x_{*},\theta)\|_{\ell_{2}(\mathcal{Y})}^{2}=0,$ (8)
where $\|f\|_{\ell_{2}(\mathcal{Y})}:=\sqrt{\sum_{y\in\mathcal{Y}}f(y)^{2}}$ denotes the $\ell_{2}$ norm. Then for $p_{x,1}$-a.e. $(x_{1:n},x_{*})$ and $p_{M}$-a.e. $y_{1:n}$, the $n\to\infty$ limit of (7) is equivalent to (7).
In the above, (8) is the assumed identifiability condition for the likelihood function. We expect the numerous technical restrictions to be relaxable, but refrain from a more general proof for brevity. The main conditions that enable the equivalence are that (i) $p_{M}$ can be augmented to define an exchangeable (or c.i.d.) model for $(x,y)$, and (ii) the likelihood function is identifiable.
Proof.
We will show that each of the two terms in (7) and (7) are equivalent.
Let us introduce the following notations: define $\mathcal{Z}:=\mathcal{X}\times\mathcal{Y}$ and, for any $x_{i},y_{i}$, $z_{i}:=(x_{i},y_{i})\in\mathcal{Z}$. For all $n\geq 0$, define $\bar{p}_{M}(z_{n+1}=(x,y)\mid z_{1:n})=p_{x,1}(x)p_{M}(y_{n+1}=y\mid x_{n+1}=x,x_{1:n},y_{1:n}),\bar{p}(z=(x,y)\mid\theta)=p_{x,1}(x)p(y\mid\theta,x)$. Clearly, $\bar{p}_{M}$ is equivalent to a Bayesian model with prior $\pi$ and the factorized likelihood $\bar{p}$. Thus, $z_{n+1:N+1}\sim\bar{p}_{M}(\cdot\mid z_{1:n})$ are exchangeable, and we have, for all $n,z_{1:n},N>n,y^{\prime}\in\mathcal{Y}$ and $p_{x,1}$-almost every $x_{*}$,
$\mathbb{E}_{\bar{p}_{M}}(r(y^{\prime},y_{N+1};x_{*})\mid z_{1:n},x_{N+1}=x_{*})=\mathbb{E}_{\bar{p}_{M}}(r(y^{\prime},y_{n+1};x_{*})\mid z_{1:n},x_{n+1}=x_{*}).$
By definitions, the terms above equal the first term in (7) and (7), respectively, with their outmost infimum removed. Retaking infimum proves the equivalence of the first terms.
For the second term, observe that by the boundedness of $r$ and the discreteness of $\mathcal{Y}$, for all $x_{*}\in\mathcal{X}$ the function $R_{x_{*}}(\rho):=\inf_{y^{\prime}\in\mathcal{Y}^{\prime}_{I}}\mathbb{E}_{y\sim\rho}r(y^{\prime},y;x_{*})$ is Lipschitz continuous w.r.t. the $\ell_{2}(\mathcal{Y})$ norm. [^5] Note that any distribution $\rho$ over $\mathcal{Y}$ can be identified with a probability mass function in $\ell_{2}(\mathcal{Y})$ since $\mathcal{Y}$ is discrete. Thus, for $\pi$-a.e. $\theta$ we have
$\lim_{N\to\infty}\mathbb{E}_{z_{1:N-n}\overset{i.i.d.}{\sim}\bar{p}(z\mid\theta)}|R_{x_{*}}(p_{M}(y=\cdot\mid x=x_{*},z_{1:N-n}))-R_{x_{*}}(p(y=\cdot\mid x=x_{*},\theta))|=0$
following (8). It follows by the boundedness of $R_{x_{*}}(\cdot)$ and the dominated convergence theorem that
$\phantom{=}\lim_{N\to\infty}\mathbb{E}_{\theta\sim\pi,z_{1:N-n}\sim\bar{p}(z\mid\theta)}R_{x_{*}}(p_{M}(y=\cdot\mid x=x_{*},z_{1:N-n}))=\mathbb{E}_{\theta\sim\pi}R_{x_{*}}(p(y=\cdot\mid x=x_{*},\theta)).$ (9)
The RHS above equals the expectation of the second term in (7) over $z_{1:n}\sim\bar{p}_{M}$. Denote the second term in (7) as $u_{N}(z_{1:n})$. By definition of $\bar{p}_{M}$ and its exchangeability, as well as the convexity of $R_{x_{*}}(\cdot)$, we have
$\mathbb{E}_{\bar{p}_{M}(z_{1:n})}u_{N}(z_{1:n})$ $=\mathbb{E}_{\bar{p}_{M}(z_{1:n})}\mathbb{E}_{\bar{p}_{M}(z_{n+1:N}\mid z_{1:n})}R_{x_{*}}(\bar{p}_{M}(y_{N+1}=\cdot\mid x_{N+1}=x_{*},z_{1:N}))$
$\leq\mathbb{E}_{\bar{p}_{M}(z_{n+1:N})}R_{x_{*}}(\mathbb{E}_{\bar{p}_{M}(z_{1:n}\mid z_{n+1:N})}\bar{p}_{M}(y_{N+1}=\cdot\mid x_{N+1}=x_{*},z_{1:N}))$
$=\mathbb{E}_{\bar{p}_{M}(z_{1:N-n})}R_{x_{*}}(\bar{p}_{M}(y_{N-n+1}=\cdot\mid x_{N-n+1}=x_{*},z_{1:N-n}))$
$=\mathbb{E}_{\theta\sim\pi,z_{1:N-n}\sim\bar{p}(z\mid\theta)}R_{x_{*}}(p_{M}(y=\cdot\mid x=x_{*},z_{1:N-n})).$
Combining with (9) and applying the dominated convergence theorem yield
$\mathbb{E}_{\bar{p}_{M}(z_{1:n})}\lim_{N\to\infty}u_{N}(z_{1:n})=\lim_{N\to\infty}\mathbb{E}_{\bar{p}_{M}(z_{1:n})}u_{N}(z_{1:n})\leq\mathbb{E}_{\bar{p}_{M}(z_{1:n})}\mathbb{E}_{\pi(\theta\mid z_{1:n})}R_{x_{*}}(p(y=\cdot\mid x=x_{*},\theta)).$ (10)
On the other hand, we have
$u_{N}(z_{1:n})$ $=\mathbb{E}_{\bar{p}_{M}(z_{n+1:N}\mid z_{1:n})}R_{x_{*}}(\bar{p}_{M}(y=\cdot\mid x=x_{*},z_{1:N}))$
$=\mathbb{E}_{\bar{p}_{M}(z_{n+1:N}\mid z_{1:n})}R_{x_{*}}(\mathbb{E}_{\pi(\theta\mid z_{1:N})}p(y=\cdot\mid x=x_{*},\theta))$
$\geq\mathbb{E}_{\bar{p}_{M}(z_{n+1:N}\mid z_{1:n})}\mathbb{E}_{\pi(\theta\mid z_{1:N})}R_{x_{*}}(p(y=\cdot\mid x=x_{*},\theta))$
$=\mathbb{E}_{\pi(\theta\mid z_{1:n})}R_{x_{*}}(p(y=\cdot\mid x=x_{*},\theta)),$
so the same holds for $\lim_{N\to\infty}u_{N}(z_{1:n})$. Comparing with (10) we find that equality must hold for $\bar{p}_{M}$-a.e. $z_{1:n}$. This proves the equivalence for the second terms in (7) and (7), and consequently the original claim. ∎
Appendix B Additional Experimental Details
In both experiments we use the following versions of LM APIs: gpt-3.5-turbo-0125, gemini-1.5-flash-001, gemini-1.5-pro-001.
The machine translation experiment is implemented as follows: for each test query $x_{*}$, we sample the ICL demonstrations $(x_{1:n},y_{1:n})$, a single set of $x_{n+1:N}$, and $5$ sets of $y_{n+1:N}\sim p_{M}(x_{1:N},y_{1:n})$; then for each path sample $y_{n+1:N}$ we draw $8$ samples from $p_{M}(y_{*}\mid x_{1:N},y_{1:N},x_{*})$. The action space is defined using $5\times 1$ samples from each path sample, and the expectations in (2) and (7) using the remaining $5\times 7$ samples. For each $x_{*}$, we reuse the same set of samples for $(x_{n+1:N},x_{1:n},y_{1:n})\mid x_{*}$ across all experiments. The completion inputs $x_{n+1:N}$ are generated using Gemini-1.5-Flash with the following prompt:
Please generate {K} sentences that convey a broadly similar meaning as the one provided, ensuring they are of similar length and use varied wording. Place a blank line after each sentence. Avoid adding any extra explanations.
{x_*}
In the above, $x_{*}$ denotes the test query. For simplicity, we report the chrF scores evaluated using the final samples $y_{*}\mid x_{1:N},y_{1:N},x_{*}$, which is similar to the reinforced ICL method (Agarwal et al.,, 2024).
Figure 3: Machine translation: reliability diagrams for a subset of experiments. Error bar indicates 95% confidence interval for the expected observed utility.

### v2

@@
 We further derive a measure for epistemic uncertainty, based on a missing data perspective and its characterization as an excess risk.
-The proposed measures can be applied to black-box language models.
-We demonstrate the proposed methods on question answering and machine translation tasks, where they extract broadly meaningful uncertainty estimates from GPT and Gemini models and quantify their calibration.
+The proposed methods can be applied to black-box language models.
+We illustrate the methods on question answering and machine translation tasks.
+Our experiments provide a principled evaluation of task-specific calibration, and demonstrate that epistemic uncertainty offers a promising deferral strategy for efficient data acquisition in in-context learning.
 1 Introduction
-We are interested in uncertainty quantification for language models (LMs) in free-form natural language generation (NLG): given instruction $I$, the model generates a response $y^{\prime}$ based on its predictive distribution $p_{M}(y\mid I)$, where $y^{\prime}$ could be any natural language passage that fits the instruction.
-An example of this is question answering (QA): given a question from the user, the model may provide a brief answer, but it may also follow with supporting facts and explanations, which can vary in form and detail.
+We are interested in uncertainty quantification (UQ) for language models (LMs) in free-form natural language generation (NLG): given instruction $I$, the model generates a response $y^{\prime}$ based on its predictive distribution $p_{M}(y\mid I)$, where $y^{\prime}$ could be any natural language passage that fits the instruction.
+An example of this is in question answering (QA): given a question from the user, the model may provide a brief answer, but it may also follow with supporting facts and explanations, which can vary in form and detail.
 The user can be satisfied by a wide variety of responses, irrespective of their style or (to some extent) the choice of supporting facts included.
@@
 This paper is about the observation that the above challenges can be resolved in a more general setup from a perspective of Bayesian decision theory (Savage,, 1954).
-We first assume that we can compute a similarity-type measure $S(y^{\prime},y;I)$ that characterizes our utility function when the model generates $y^{\prime}$ in respond to $I$ and $y$ is a (true or hypothetical) correct response.
+We start by assuming that we can compute a similarity-type measure $S(y^{\prime},y;I)$ that characterizes our utility function when the model generates $y^{\prime}$ in respond to $I$ and $y$ is a (true or hypothetical) correct response.
 This generalizes previous works based on semantic equivalence, which corresponds to defining $S$ using entailment.
-It can also be applied to traditional structured prediction tasks where a choice of $S$ is readily available (e.g., the chrF score for machine translation, Popović,, 2015), as well as tasks where evaluation can be adequately implemented using frontier language models that computes $S$ based on few-shot demonstrations or detailed instructions.
+It can also be applied to traditional structured prediction tasks where a choice of $S$ is readily available (e.g., lexical scores for machine translation, Koehn,, 2009), as well as any task where evaluation can be implemented using LMs that compute $S$ based on few-shot demonstrations or detailed instructions (e.g., Yu et al.,, 2023; Wang et al.,, 2023; Bannur et al.,, 2024).
 To quantify the subjective uncertainty defined by the LM, we assume the generation $y$ is always chosen to maximize the expected utility under the LM’s predictive distribution $p_{M}$2.1).
-This is a common assumption in NLG (Bickel and Doksum,, 1997; Bertsch et al.,, 2023).
-It can be supported by a common belief that high-capacity LMs may approximate Bayesian inference (e.g., Xie et al.,, 2021; Akyürek et al.,, 2022; Hahn and Goyal,, 2023; Ye et al.,, 2024), and is also more broadly applicable whenever we view $p_{M}$ as our best available model for $y\mid I$.
+This is a common, albeit implicit, assumption in NLG (Bickel and Doksum,, 1997; Bertsch et al.,, 2023).
+It can be supported by a belief that high-capacity LMs may approximate Bayesian inference (e.g., Xie et al.,, 2021; Akyürek et al.,, 2022; Hahn and Goyal,, 2023), and is also more broadly applicable whenever we view $p_{M}$ as our best available model for $y\mid I$.
 Subjective uncertainty is then naturally characterized by the Bayes risk, or equivalently the maximum achievable expected utility, where the action space is defined by the candidate generations available.
 This simple observation allows us to understand that previous methods for “semantic uncertainty” can be adapted to a broader range of scenarios, regardless of whether semantic equivalence is relevant.
-It also provides a unique, principled approach for aggregating similarity measures among generations (§2.2).
-For subjective uncertainty measures to be useful, it will be ideal if the LM is calibrated.
-It has been unclear in previous works how calibration can be evaluated in free-form NLG tasks, given the distinction between relevant and irrelevant differences for a specific task and our apparent inability to formally characterise it.
-The decision-theoretic view provides a natural answer to this question: an LM is deemed as to possess a calibrated notion of uncertainty if the expected subjective utility of its actions matches the expectation of the actually incurred utility, defined w.r.t.
-responses from the true data distribution (§2.3).
-This observation allows us to quantify the calibration of LMs through reliability diagrams (Murphy and Winkler,, 1977) and a generalized version of expected calibration error (Naeini et al.,, 2015).
-In the discussion we also note a generally applicable definition for “task-specific semantics”: the equivalence class defined by the action-utility map, known as Savage acts.
-Bayesian modelling allows for the decomposition of predictive uncertainty into epistemic uncertainty and aleatoric uncertainty (Der Kiureghian and Ditlevsen,, 2009).
-The decomposition allows us to better understand and diagnose the model’s uncertainty (e.g., only the epistemic uncertainty should decrease as the training set grows); it can also guide data acquisition (Kendall and Gal,, 2017).
-The quantification of epistemic uncertainty in general NLG applications appears challenging, due to the black-box nature of LMs and the free-form generations involved.
-§2.4 demonstrates that the decision-theoretic view enables us to quantify epistemic uncertainty for in-context learning (ICL, Brown et al.,, 2020) applications; key additional ingredients include a missing data perspective to Bayesian modelling (Fong et al.,, 2024) and a connection between epistemic uncertainty and a notion of excess risk (Xu and Raginsky,, 2022).
-The resulted measure is connected to and generalizes various previous works on epistemic uncertainty quantification for black-box models, which provides independent justification for its use.
-The rest of this paper is structured as follows: §2 presents the proposed methodology and discusses connections to previous work; §3 provides experimental illustrations; and §4 provides concluding remarks.
-2 Quantifying Subjective Uncertainty and Calibration
+It also highlights a unique, principled approach for aggregating similarity measures among generations (§2.2).
+For subjective uncertainty measures to be useful, it will be ideal if the LM is more calibrated.
+It has been unclear in previous works how calibration can be evaluated in general NLG tasks, given the distinction between relevant and irrelevant differences for a task and our apparent inability to formally characterise it.
+The decision-theoretic view provides a natural answer to this question: an LM is deemed to possess a calibrated notion of uncertainty if the expected subjective utility of its action matches the expectation of the actually incurred utility w.r.t.
+the true data distribution (§2.3).
+This observation allows us to quantify the calibration of LMs through reliability diagrams (Murphy and Winkler,, 1977) and a generalization of the expected calibration error (Naeini et al.,, 2015).
+Bayesian modelling enables the decomposition of predictive uncertainty into epistemic uncertainty and aleatoric uncertainty (Der Kiureghian and Ditlevsen,, 2009).
+Epistemic uncertainty refers to the reducible proportion of uncertainty and can guide data acquisition (Kendall and Gal,, 2017), or more broadly, indicate when additional information can be most effectively used to improve prediction.
+The quantification of epistemic uncertainty is of interest in many LM applications, as predictions can often benefit from additional information (Lewis et al.,, 2020; Brown et al.,, 2020; Li et al.,, 2023), which may be costly to obtain (Agarwal et al.,, 2024; Min et al.,, 2024).
+Yet it appears challenging due to the black-box nature of LMs and the free-form generations involved.
+As we show in §2.4, the decision-theoretic view leads to a principled measure of task-specific epistemic uncertainty in in-context learning (ICL, Brown et al.,, 2020); key additional ingredients include a missing data perspective to Bayesian modelling (Fong et al.,, 2023) and a connection between epistemic uncertainty and an excess risk (Xu and Raginsky,, 2022).
+The resulted measure is connected to and generalizes several recent methods, which provides independent justification for its use.
+The methodological developments allow us to better understand the “probabilistic uncertainty” in modern LMs, that is, the uncertainty reflected in the variation of the model’s predictive distribution.
+§4 provides experimental illustrations: We revisit past comparisons (Tian et al.,, 2023; Lin et al.,, 2024) of probabilistic and non-probabilistic UQ for instruction-tuned LMs by conducting a principled evaluation of task-specific calibration (§4.1).
+In §4.2, we demonstrate that epistemic uncertainty provides a deferral strategy (Madras et al.,, 2018; Dohan et al.,, 2022) that facilitates efficient data acquisition in many-shot ICL (Agarwal et al.,, 2024).
+The rest of this paper is structured as follows: §2 presents the proposed methodology, §3 reviews related work, §4 presents the experiments, and §5 provides concluding remarks.
+2 Quantifying Uncertainty and Calibration
 2.1 A Utilitarian Setup
-The basic setup of this paper is as follows:
+The setup of this paper is as follows:
   • we have a NLG task: given prompt $I\in\mathcal{I}$, generate a response $y^{\prime}\in\mathcal{Y}$ where $\mathcal{Y}$ denotes the space of natural language responses;
-  • our utility can be measured through the similarity measure $S(y^{\prime},y;I)\in\mathbb{R}$, which can be cheaply evaluated;
-  • we have access to an LM $p_{M}$, and, in the absence of further evidence, consider $p_{M}$ sufficiently trustworthy so that the ideal generation (i.e., action) $y^{\prime}$ given $I$ should maximize the expected utility,
+  • our utility can be quantified through a similarity measure $S(y^{\prime},y;I)\in\mathbb{R}$, which can be cheaply evaluated;
+  • we have an LM with predictive distribution $p_{M}$, and, in the absence of further evidence, consider $p_{M}$ sufficiently trustworthy so that the ideal generation (i.e., action) $y^{\prime}$ for $I$ should maximize the expected utility,
     $y^{\prime}:=\operatorname*{arg\,max}_{y^{\prime}\in\mathcal{Y}^{\prime}_{I}}\mathbb{E}_{y\sim p_{M}(\cdot\mid I)}S(y^{\prime},y;I),$ (1)
-    where $\mathcal{Y}^{\prime}_{I}\subset\mathcal{Y}$ denotes the space of candidate generations.
+    where $\mathcal{Y}^{\prime}_{I}\subset\mathcal{Y}$ is the space of candidate generations.
 In the above, $p_{M}$ represents a subjective belief about the true data distribution.
-$\mathcal{Y}_{I}^{\prime}\subset\mathcal{Y}$ can be determined based on computational constraints, e.g., as a separate set of samples from $p_{M}(\cdot\mid I)$.111 When $\mathcal{Y}_{I}^{\prime}$ is stochastic, expected utility/risk quantities such as (1) should also be averaged over the randomness in $\mathcal{Y}_{I}^{\prime}$.
-For brevity, we omit such (outmost) expectations throughout the paper.
+$\mathcal{Y}_{I}^{\prime}\subset\mathcal{Y}$ can be determined based on computational constraints, e.g., as a separate set of samples from $p_{M}(\cdot\mid I)$.111 When $\mathcal{Y}_{I}^{\prime}$ is stochastic, expectation quantities such as (1) should also be averaged over the randomness in $\mathcal{Y}_{I}^{\prime}$.
+For brevity, we will omit all such (outmost) expectations.
 The utility $S$ determines a risk function $r(y^{\prime},y;I)=-S(y^{\prime},y;I)$, which can be scaled and shifted as desired.
@@
 This decision-theoretic perspective is taken in the framework of minimum Bayes risk (MBR) decoding (Bickel and Doksum,, 1997), which advocates for generating $y$ based precisely on (1).
-MBR recovers many modern methods for natural language generation (Bertsch et al.,, 2023), and these methods can be viewed as having adopted the same view implicitly.
-Note that adopting the decision-theoretic setup does not compel us to assume that $p_{M}$ is a strictly Bayesian model (e.g., in the sense of Xie et al.,, 2021): although that will provide a good justification, it is also possible that a non-Bayesian LM may better approximate our belief than any available Bayesian model does.
-2.2 Task-Specific Measure of Subjective Uncertainty
+MBR recovers many modern methods for natural language generation (Bertsch et al.,, 2023), which can be viewed as implicitly adopting the same view.
+Note that the decision-theoretic setup does not compel us to assume that $p_{M}$ is strictly a Bayesian model (e.g., in the sense of Xie et al.,, 2021): although that will provide a good justification, it is also possible that a non-Bayesian LM better approximates our belief than any available Bayesian model does.
+2.2 Task-Specific Measures of Subjective Uncertainty
 It is well known that Bayesian uncertainty measures are connected to minimum achievable risks of the form
-$R_{B}(I;p_{M},\mathcal{Y}^{\prime}_{I}):=\min_{y^{\prime}\in\mathcal{Y}^{\prime}_{I}}\mathbb{E}_{y\sim p_{M}(\cdot\mid I)}r(y^{\prime},y;I).$ (2)
-For example, if $\mathcal{Y}=\mathcal{Y}^{\prime}_{I}\subset\mathbb{R}$ and $r(y,y^{\prime};I)=(y-y^{\prime})^{2}$, $R_{B}(I;p_{M},\mathcal{Y}^{\prime}_{I})$ will be equivalent to the variance of the predictive distribution; if with an abuse of notation we redefine $y^{\prime}$ as a probability density function and $r(y^{\prime},y;I)\leftarrow\log y^{\prime}(y)$, $R_{B,\mathcal{Y}^{\prime}_{I}}(I)=H[p_{M}(\cdot\mid I)]$ will be the entropy.
-When the action space $\mathcal{Y}^{\prime}_{I}\subsetneq\mathcal{Y}$ is a strict subset, (2) will be connected to measures of usable information as discussed in Xu et al., (2019).
-It is thus natural to also adopt (2) as a measure of subjective uncertainty for NLG tasks. To implement (2), we adopt Monte Carlo approximation for the expectation.
-A common generation strategy is to follow the Gibbs predictor, which corresponds to (1) with $\mathcal{Y}^{\prime}_{I}:=\{y_{I}\}$ and $y_{I}\sim p_{M}(\cdot\mid I)$ being a single sample.
-In such cases, the expected uncertainty $\mathbb{E}_{\mathcal{Y}^{\prime}_{I}}R_{B}(I;p_{M},\mathcal{Y}^{\prime}_{I})$ is connected to the degree statistic in Lin et al., (2024, Eq.
-8); the difference is that we use a similarity measure that is assumed to define our utility.
-While Lin et al., (2024) explored 9 different uncertainty measures for the Gibbs predictor, the decision-theoretic perspective leads to one unique method without any underspecified design choices, justifies its use beyond the scope of semantic equivalence, and highlights that different choices of predictors ($\mathcal{Y}_{I}^{\prime}$) should be matched with different uncertainty measures.
-However, being a measure of subjective uncertainty, it is not guaranteed to always improve on downstream tasks compared with heuristic alternatives, especially if the LM is less calibrated.
+$R_{B}(I):=\min_{y^{\prime}\in\mathcal{Y}^{\prime}_{I}}\mathbb{E}_{y\sim p_{M}(\cdot\mid I)}r(y^{\prime},y;I).$ (2)
+For example, if $\mathcal{Y}=\mathcal{Y}^{\prime}_{I}\subset\mathbb{R}$ and $r(y,y^{\prime};I)=(y-y^{\prime})^{2}$, $R_{B}(I)$ will be equivalent to the predictive variance; if with an abuse of notation we redefine $y^{\prime}$ as a density function and $r(y^{\prime},y;I)\leftarrow\log y^{\prime}(y)$, $R_{B,\mathcal{Y}^{\prime}_{I}}(I)$ will become the entropy.
+When the action space $\mathcal{Y}^{\prime}_{I}\subsetneq\mathcal{Y}$ is a strict subset, (2) becomes connected to measures of “usable information” discussed in Xu et al., (2019).
+It is thus natural to also adopt (2) as a measure of subjective uncertainty in NLG.
+By definitions, it represents the task-specific uncertainty we are interested in, and we recover measures of semantic uncertainty if we define $r$ based on semantic equivalence.
+(2) can be implemented through Monte-Carlo estimation.
+A common generation strategy is defined by the Gibbs predictor, which corresponds to (1) with $\mathcal{Y}^{\prime}_{I}:=\{y_{I}\}$, where $y_{I}\sim p_{M}(\cdot\mid I)$ is a single random sample.
+In such cases, the expected uncertainty $\mathbb{E}_{\mathcal{Y}^{\prime}_{I}}R_{B}(I)$ is connected to the degree statistic in Lin et al., (2024, Eq.
+8); see also Wang et al., (2022, p.8).
+The difference is that we use a similarity measure that we assume to define our utility.
+Compared with previous works (Kuhn et al.,, 2023; Lin et al.,, 2024; Nikitin et al.,, 2024) which explored many uncertainty measures based on the aggregation of similarity measures, the decision-theoretic perspective highlights one unique method without any underspecified design choices, justifies its use beyond the scope of semantic equivalence, and shows that different choices of predictors ($\mathcal{Y}_{I}^{\prime}$) should be matched with different uncertainty measures.
+Although being a measure of subjective uncertainty, it is not guaranteed to always outperform heuristic alternatives on downstream tasks, in particular if the LM is severely miscalibrated.
 2.3 Evaluation of Task-Specific Calibration
-Subjective uncertainty measures are most useful when the LM is calibrated.
-As discussed in Huang et al., (2024), many previous works studying uncertainty on NLG tasks did not distinguish between calibration and predictive performance in their evaluation, which leads to a relative lack of understanding about the calibration of LMs in such scenarios.
-A typical definition of calibration is as follows (see e.g., Johnson et al.,, 2024): suppose the spaces $\mathcal{Y},\mathcal{I}$ are discrete, and denote the true data distribution as $p_{0}$; a model $p_{M}$ is considered calibrated w.r.t.
+Subjective uncertainty measures are most useful when the LM is reasonably calibrated.
+As discussed in Huang et al., 2024a , many previous works studying uncertainty in free-form NLG did not distinguish between calibration and predictive performance in their evaluation, which leads to a relative lack of understanding about the calibration in such settings.
+A common definition of calibration is as follows (see e.g., Johnson et al.,, 2024): suppose the spaces $\mathcal{Y},\mathcal{I}$ are discrete, and denote the true data distribution as $p_{0}$; a model $p_{M}$ is considered calibrated w.r.t.
 some grouping function $G(I):\mathcal{I}\to\mathcal{G}$ if we have
 $\mathbb{E}_{p_{0}}(p_{M}(y=y^{\prime}\mid I)\mid G(I)=g)=\mathbb{E}_{p_{0}}(p_{0}(y=y^{\prime}\mid I)\mid G(I)=g)~{}~{}\forall g\in\mathcal{G},y^{\prime}\in\mathcal{Y}.$ (3)
-The grouping function $G$ may be defined based on the model’s prediction, or it may incorporate additional information such as input demographics.
-In the NLG setting, however, the criterion (3) can be trivially violated if, e.g., the LM always generates a longer response compared with $p_{0}$.
-In light of the assumptions in §2.1, it is natural to relax (3) as follows:
-$\mathbb{E}(p_{M}(\{y:S(y^{\prime},y;\mathcal{I})=s\}\mid I)\mid G(I)=g)$ $=\mathbb{E}(p_{0}(\{y:S(y^{\prime},y;\mathcal{I})=s\}\mid I)\mid G(I)=g)$
-$\hskip 130.0002pt~{}~{}\forall g\in\mathcal{G},y^{\prime}\in\mathcal{Y}^{\prime}_{I},s\in\mathbb{R}.$ (3’)
+$G(I)$ can be defined based on uncertainty measures, or it may incorporate additional information such as input demographics.
+In NLG, however, (3) can be trivially violated if, e.g., the LM always generates a more verbose response compared with $p_{0}$, which may happen due to instruction tuning (Adlakha et al.,, 2023).
+In light of the assumptions in §2.1, it is natural to relax (3) to require that for all $g\in\mathcal{G},y^{\prime}\in\mathcal{Y}^{\prime}_{I},s\in\mathbb{R}$
+$\mathbb{E}(p_{M}(\{y:S(y^{\prime},y;\mathcal{I})=s\}\mid I)\mid G(I)=g)$ $=\mathbb{E}(p_{0}(\{y:S(y^{\prime},y;\mathcal{I})=s\}\mid I)\mid G(I)=g)$ (3’)
 (3) is equivalent to (3) if $\mathcal{Y}^{\prime}_{I}=\mathcal{Y}$ and $S(y^{\prime},y;\mathcal{I})=\mathbf{1}\{y=y^{\prime}\}$, or more generally if for all $I\in\mathcal{I}$ the map from $y$ to the function
-$s_{y,I}:\mathcal{Y}_{I}^{\prime}\to\mathbb{R},~{}y^{\prime}\mapsto S(y^{\prime},y;I)$ (4)
-is injective (e.g., if $S$ is defined by a metric on $\mathcal{Y}$).
-(3) becomes weaker if $S(y^{\prime},y;\mathcal{I})$ only depends on $y$ and $y^{\prime}$ through a non-invertible function, e.g., a map from a natural language passage to its semantics; in such cases, it is clear that (3) remains a sufficiently strong criterion for task-specific calibration.
-More generally, we can always view the function $s_{y,I}(y^{\prime})$ as a representation of “task-specific semantics”.222 For the goals of providing the best generation from $\mathcal{Y}^{\prime}_{I}$ and quantifying its risk, it suffices to define the function $s_{y,I}$ on $\mathcal{Y}^{\prime}_{I}$.
-For purposes beyond the scope of the generation task, it may be helpful to extend its domain to a larger subset of $\mathcal{Y}$.
-In fact, the replacement of $y^{\prime}\in\mathcal{Y}^{\prime}_{I}$ with the more compact $s_{y^{\prime},I}(\cdot)$ has been advocated by Savage, (1954), and $s_{y,I}$ is referred to as a Savage act (see e.g., Marinacci,, 2015, §2.4).
-A main reason we are interested in criteria such as (3) or (3) is that they guarantee the calibration of uncertainty measures.
+$s_{y,I}:\mathcal{Y}_{I}^{\prime}\to\mathbb{R},~{}y^{\prime}\mapsto S(y^{\prime},y;I)$
+is injective.
+(3) becomes weaker if $S$ only depends on $y$ and $y^{\prime}$ through a non-invertible function, e.g., a map from a natural language passage to its semantics.
+In such cases, it is clear that (3) remains sufficiently strong while focusing on task-specific calibration.
+Remark 2.1.
+We can generally view the function $s_{y,I}(y^{\prime})$ as a representation of the “task-specific semantics”.222 For the purposes of MBR generation and uncertainty quantification, it suffices to define the function $s_{y,I}$ on $\mathcal{Y}^{\prime}_{I}$.
+For other purposes it may be helpful to extend its domain.
+In fact, the replacement of $y^{\prime}\in\mathcal{Y}^{\prime}_{I}$ with $s_{y^{\prime},I}(\cdot)$ has been advocated by Savage, (1954), and $s_{y,I}$ is referred to as a Savage act (see e.g., Marinacci,, 2015, §2.4).
+A main reason we are interested in (3) or (3) is that they guarantee the calibration of uncertainty measures.
 It follows from (3) that $\mathbb{E}_{p_{0}(I\mid G(I)=g)p_{M}(y\mid I)}r(y^{\prime},y;I)=\mathbb{E}_{p_{0}(I\mid G(I)=g)p_{0}(y\mid I)}r(y^{\prime},y;I)$ for all $y^{\prime}$, and thus
-$\mathbb{E}_{p_{0}(I\mid G(I)=g)}R_{B}(I;p_{M},\mathcal{Y}^{\prime}_{I})=\mathbb{E}_{p_{0}(I\mid G(I)=g)p_{M}(y\mid I)}r(\hat{y}_{I},y;I)=\mathbb{E}_{p_{0}(I\mid G(I)=g)p_{0}(y\mid I)}r(\hat{y}_{I},y;I),$ (5)
-where $\hat{y}_{I}$ denotes the MBR generation (1). Plugging in $G(I)=R_{B}(I;p_{M},\mathcal{Y}^{\prime}_{I})$, we find that LMs are calibrated in the sense of (3) only if
-$\text{gECE}(p_{M}):=\mathbb{E}_{s}|f_{M}(s)-s|=0,~{}~{}~{}\text{where }f_{M}(s):=\mathbb{E}_{I,y,\hat{y}_{I}}(r(\hat{y}_{I},y;I)\mid R_{B}(I;p_{M},\mathcal{Y}^{\prime}_{I})=s)~{}\forall s\in\mathbb{R}.$ (6)
-(6) provides a natural generalization of the expected calibration error (ECE, Naeini et al.,, 2015) in classification, which it recovers with $r(y^{\prime},y;I)\leftarrow\mathbf{1}\{y^{\prime}\neq y\}$.
-Similar to the classification case, (6) can be estimated through histogram binning.
-The plot for the function $f_{M}$ is called a reliability diagram (Murphy and Winkler,, 1977) and provides information on overconfidence ($f_{M}(s)<s$) or underconfidence.
-Importantly, these methods remain applicable in NLG as we have shown.
-Moreover, they will automatically focus on the task-specific aspects of uncertainty, since the function $f_{M}$ only depends on model generations through the utility $S$.
-Remark 2.1.
-We discussed the evaluation of LM calibration based on the uncertainty measure (2).
-To our knowledge, only Huang et al., (2024) studied a similar problem in the general NLG setting, namely the rank-calibration of general uncertainty measures.
-As discussed in Huang et al., (2024, §4.2), rank-calibration is still a different property; in particular, (2) will be rank-calibrated whenever $f_{M}$ is monotone, even though it can be arbitrarily overconfident or underconfident and thus miscalibrated.
+$\mathbb{E}_{p_{0}(I\mid G(I)=g)}R_{B}(I)=\mathbb{E}_{p_{0}(I\mid G(I)=g)p_{0}(y\mid I)}r(\hat{y}_{I},y;I),$
+where $\hat{y}_{I}$ denotes the MBR generation (1). Plugging in $G(I)=R_{B}(I)$, we find that the calibration criterion (3) holds only if
+$\text{ECE}(p_{M}):=\mathbb{E}_{s}|f_{M}(s)-s|=0,~{}~{}~{}\text{where }f_{M}(s):=\mathbb{E}_{I,y,\hat{y}_{I}}(r(\hat{y}_{I},y;I)\mid R_{B}(I)=s)~{}\forall s\in\mathbb{R}.$ (4)
+(4) generalizes the expected calibration error (ECE, Naeini et al.,, 2015) in classification, which it recovers with $r(y^{\prime},y;I)\leftarrow\mathbf{1}\{y^{\prime}\neq y\}$.
+As in the classification case, we can estimate (4) through histogram binning.
+The plot for $f_{M}$ is called a reliability diagram (Murphy and Winkler,, 1977) and provides insights on overconfidence ($f_{M}(s)<s$) or underconfidence.
+Importantly, these methods remain applicable in NLG without the need to introduce any binary notion of correctness.
+Moreover, they will automatically focus on the task-specific aspects of generation, as ECE and $f_{M}$ only depend on model samples through the utility.
 2.4 Representing and Eliciting Epistemic Uncertainty
-We now turn to the problem of epistemic uncertainty quantification in in-context learning, where the prompt $I=\langle x_{1},y_{1},\ldots,x_{n},y_{n},x_{*}\rangle$ is formed by $n$ demonstrations $(x_{i},y_{i})$ followed by a test input $x_{*}$.
-We assume that $r(y,y^{\prime};I)=r(y,y^{\prime};x_{*})$ only depends on $I$ through $x_{*}$, which is commonly satisfied in applications.
-We propose the following measure of epistemic uncertainty inspired by Fong et al., (2024); Xu and Raginsky, (2022):
-$\inf_{y^{\prime}\in\mathcal{Y}^{\prime}_{I}}\mathbb{E}_{p_{x}(x_{n+1:N})}\mathbb{E}_{p_{M}(y_{n+1:N}\mid y_{1:n},x_{1:N})}\mathbb{E}_{p_{M}(y_{*}\mid y_{1:N},x_{1:N},x_{*})}r(y^{\prime},y_{*};x_{*})$
-$\hskip 40.00006pt-\mathbb{E}_{p_{x}(x_{n+1:N})}\mathbb{E}_{p_{M}(y_{n+1:N}\mid y_{1:n},x_{1:N})}\inf_{y^{\prime}\in\mathcal{Y}^{\prime}_{I}}\mathbb{E}_{p_{M}(y_{*}\mid y_{1:N},x_{1:N},x_{*})}r(y^{\prime},y_{*};x_{*}).$ (7)
-In the above, $p_{x}$ denotes a user-specified distribution of additional ICL inputs, $y_{n+1:N}$ are model generations, $N>n$ is a hyperparameter, and we assume the LM has good performance for ICL up to a sample size of $N$.
-Eq.
-(7) can be estimated through autoregressive sampling.
-As long as we use the same set of samples for its both terms, the resulted estimate will always be non-negative, due to the convexity of the functional $\rho\mapsto\inf_{y^{\prime}\in\mathcal{Y}^{\prime}_{I}}\mathbb{E}_{\rho(y_{*})}r(y^{\prime},y_{*};x_{*})$.
-To understand (7), first consider a simplified scenario as follows: suppose the LM $p_{M}$ is equivalent to a Bayesian model that assumes $\{(x_{i},y_{i})\}$ are i.i.d.
-conditional on a latent variable $\theta$, $p_{x}$ defines i.i.d.
-$x_{n+1:N}$, and $\{x_{n+i},y_{n+i}\}$ are sufficiently informative so that the likelihood function $p(y=\cdot\mid x=x_{*},\theta)$ determined by $\theta\mid x_{1:N},y_{1:N}$ is identifiable given infinite samples.
+We now turn to the quantification of epistemic uncertainty, using ICL as a key example.
+Epistemic uncertainty is the proportion of uncertainty that can be reduced given additional observations.
+It is particularly relevant when such observations are costly to obtain or incorporate into the generation process.
+In these scenarios, we can limit their use to queries with a higher level of reducible (i.e., epistemic) uncertainty.
+Epistemic uncertainty in ICL.
+In ICL the LM is prompted with $I=\langle z_{1:n},x_{*}\rangle=\langle x_{1},y_{1},\ldots,x_{n},y_{n},x_{*}\rangle$, where $z_{i}:=(x_{i},y_{i})$ are demonstration input-output pairs and $x_{*}$ denotes the test query.
+$n$ can be arbitrary.
+We assume that the user’s utility does not depend on $z_{1:n}$ given $x_{*}$, so that the risk has the form $r(y,y^{\prime};I)\equiv r(y,y^{\prime};x_{*})$.
+We propose the following measure of epistemic uncertainty, inspired by Fong et al., (2023); Xu and Raginsky, (2022):
+$\inf_{y^{\prime}\in\mathcal{Y}^{\prime}_{I}}\mathbb{E}_{p_{x}(x_{n+1:N}\mid I)}\mathbb{E}_{p_{M}(y_{n+1:N}\mid y_{1:n},x_{1:N})}\mathbb{E}_{p_{M}(y_{*}\mid z_{1:N},x_{*})}r(y^{\prime},y_{*};x_{*})$
+$\hskip 40.00006pt-\mathbb{E}_{p_{x}(x_{n+1:N}\mid I)}\mathbb{E}_{p_{M}(y_{n+1:N}\mid y_{1:n},x_{1:N})}\inf_{y^{\prime}\in\mathcal{Y}^{\prime}_{I}}\mathbb{E}_{p_{M}(y_{*}\mid z_{1:N},x_{*})}r(y^{\prime},y_{*};x_{*}).$ (5)
+In the above, $p_{x}$ denotes a user-specified distribution of inputs, $y_{n+1:N}\sim\prod_{j=n+1}^{N}p_{M}(y_{j}\mid z_{1:j-1},x_{j})$ follow the LM’s predictive distribution, and $N>n$ is a hyperparameter.
+(5) can be estimated through autoregressive sampling and Monte-Carlo estimation; see §4.2 for implementation details.
+Comparing with (2), we can see that (5) is the difference between the total uncertainty from two predictors; the second predictor has access to the additional demonstrations $z_{n+1:N}=(x_{n+1:N},y_{n+1:N})$, where $y_{n+1:N}$ are model generations.
+Thus, (5) is a subjective measure of the reducible uncertainty given $x_{n+1:N}$.
+By convexity of the $\inf$ functional, (5) is always non-negative; so is its Monte-Carlo estimate as long as we use the same set of samples for its both terms.
+Bayesian justifications.
+To further understand (5), consider a simplified setting as follows: suppose $p_{M}$ is equivalent to a Bayesian model that assumes $z_{1:n}$ are i.i.d.
+conditional on a latent variable $\theta$, $x_{n+1:N}\sim p_{x}(\cdot\mid I)$ are conditionally i.i.d., and $z_{n+1:N}$ are sufficiently informative so that the likelihood function $p(y=\cdot\mid x=x_{*},\theta)$ determined by $\theta\mid z_{1:N}$ is identifiable given infinite samples.
 Then under mild technical conditions, as $N\to\infty$ (5) will become equivalent to
-$\min_{y^{\prime}\in\mathcal{Y}^{\prime}_{I}}\mathbb{E}_{p_{M}(\theta\mid x_{1:n},y_{1:n})}\mathbb{E}_{p(y\mid\theta,x=x_{*})}r(y^{\prime},y;x_{*})-\mathbb{E}_{p_{M}(\theta\mid x_{1:n},y_{1:n})}\min_{y^{\prime}\in\mathcal{Y}^{\prime}_{I}}\mathbb{E}_{p(y\mid\theta,x=x_{*})}r(y^{\prime},y;x_{*}).$ (7’)
-The equivalence is a consequence of Doob, (1949)’s theorem and can be proved with the same idea as Fong et al., (2024).
-For completeness, we provide a proof in Appendix A.
-(7) is a subjective notion of excess risk.
-In other words, it is the amount of reducible risk should we have full knowledge about the distribution of $y\mid x=x_{*}$, which can be determined given either the latent $\theta$, or an infinite amount of missing data $\{x_{>n},y_{>n}\}$.
-As noted in Xu and Raginsky, (2022), this is precisely the intuition behind Bayesian epistemic uncertainty, and we recover standard epistemic uncertainty measures (e.g., posterior variance for a regression mean, or mutual information) if we return to simple scenarios where $r$ is the square or log loss.
-Within the Bayesian framework, (7) with a finite choice of $N$ will provide a lower bound for (7): by convexity the second term may become larger.
-The intuition is that knowledge of the samples $\{x_{n+1:N},y_{n+1:N}\}$ does not always allow for the full reduction of risk.
-It is possible to view the lower bound as an indication of the “true epistemic uncertainty” (7), and it can be tightened if we choose $\{x_{n+i}\}$ to be more similar to $x_{*}$.
-The bound will also be tight regardless of the choice of $N$ if the uncertainty is “fully aleatoric” or “fully epistemic”, i.e., if the risk for $y\mid x=x_{*}$ cannot be reduced by any number of additional observations $(x_{n+1:N},y_{n+1:N})$ or can be fully reduced using a single $(x_{n+1},y_{n+1})$.
-Alternatively, we can observe that we do not always need the tightest bound for (7): if all we can do is to collect $N-n$ real samples, the remaining proportion of the risk will be effectively irreducible.
-As discussed in §1, the Bayesian view above may be justified through a common belief that high-capacity LMs may approximate Bayesian inference.
-It is also supported by the results of Wen et al., (2022, §2.6, §4), which imply that in a multi-task setting, any $p_{M}$ with a strong average-case performance for prediction will yield a similar value for (7) as the true Bayesian model.
-Falck et al., (2024) investigated this belief empirically on synthetic low-dimensional datasets, and found it to be valid for smaller choices of $N$, although it breaks down as $N\to\infty$ becomes larger.333 Falck et al., (2024) also studied epistemic uncertainty quantification for ICL but restricted to datasets where the method of Fong et al., (2024) is applicable.
-Their method assumes $\{x_{n+1:N},$ $y_{n+1:N}\}$ can be modelled by a low-dimensional statistical model, so that a posterior for the model parameter can be constructed and visualized.
-This is unrealistic for free-form natural language data.
-Taken together, these results suggest that the Bayesian perspective could be relevant when we restrict $N$ to a scope where ICL is sample efficient.
-Independent to the Bayesian perspective, Eq.
+$\min_{y^{\prime}\in\mathcal{Y}^{\prime}_{I}}\mathbb{E}_{p_{M}(\theta\mid z_{1:n})}\mathbb{E}_{p(y\mid\theta,x=x_{*})}r(y^{\prime},y;x_{*})-\mathbb{E}_{p_{M}(\theta\mid z_{1:n})}\min_{y^{\prime}\in\mathcal{Y}^{\prime}_{I}}\mathbb{E}_{p(y\mid\theta,x=x_{*})}r(y^{\prime},y;x_{*}).$ (5’)
+The equivalence can be proved with the same idea as Fong et al., (2023); we provide a proof in App.
+A for completeness.
+(5) represents the amount of reducible risk should we have full knowledge about the distribution of $y\mid x=x_{*}$, which can be determined given either the latent $\theta$ or an infinite amount of missing data.
+As noted in Xu and Raginsky, (2022), This is precisely the intuition behind Bayesian epistemic uncertainty, and we recover standard epistemic uncertainty measures (e.g., posterior variance for a regression mean, or mutual information) if we return to the standard choices of $r$ such as the square or log loss.
+Within the Bayesian framework, (5) with a finite choice of $N$ will provide a lower bound for (5): by convexity the second term will be larger.
+The intuition is that knowledge of the samples $z_{n+1:N}$ does not always enable the full reduction of risk.
+We can view the lower bound as an indication of the “true” epistemic uncertainty (5), and it can be tightened if we choose $\{x_{n+i}\}$ to be more similar to $x_{*}$.
+The bound will also be tight for finite $N$ if the uncertainty is “fully aleatoric” or “fully epistemic”, i.e., if the risk at $x=x_{*}$ cannot be reduced by any amount of additional $z_{n+1:N}$ or if it can be fully reduced by a single $z_{n+1}$.
+Alternatively, we can observe that we do not always need the tightest bound for (5): if all we can do is to collect $N-n$ real samples, the remaining uncertainty will be effectively irreducible.
+As discussed in §1, this Bayesian view may be justified through a common belief that high-capacity LMs may approximate Bayesian inference.
+It is also supported by the results of Wen et al., (2022, §2.6, §4), which imply that in a multi-task setting, any $p_{M}$ with a strong average-case performance for prediction will yield a similar value for (5).
+Falck et al., (2024) investigated this belief empirically on synthetic numerical datasets and found it to be valid for smaller choices of $N$.
+Taken together, these results suggest that the Bayesian perspective can be relevant when we restrict $N$ to a regime where ICL is statistically efficient.
+Connection to non-Bayesian methods.
+Independent to the above, Eq.
 (5) can also be understood through its connection to recent works.
-Collier et al., (2022) studied a broadly similar problem where a proportion of the combined uncertainty can be reduced by conditioning on privilege information; this leads to a similar measure for reducible uncertainty as (7), if we view $(x_{n+1:N},y_{n+1:N})$ as the latent “privilege information”.
-Johnson et al., (2024); Ahdritz et al., (2024) proposed methods that measure the correlation between consecutive predictions $(y_{n+1},y_{*})$ given the same input $x_{n+1}=x_{*}$; as argued in Johnson et al., (2024), any learning system $p_{M}$ with complete confidence for the true data distribution should lead to conditionally i.i.d.
+Johnson et al., (2024); Ahdritz et al., (2024) proposed methods that measure the correlation between consecutive predictions $(y_{n+1},y_{*})$ for the same input $x_{n+1}=x_{*}$; as argued in Johnson et al., (2024), any learning system $p_{M}$ with complete confidence should lead to conditionally i.i.d.
 $(y_{*},y_{n+1})$ given $I$, and any $p_{M}$ with a “fully epistemic” uncertainty should produce $y_{*}=y_{n+1}$.
-It is clear that (7) matches these behaviours at the extremes: it equals 0 in the former case, and correctly indicate a full reduction of risk in the latter case (i.e., it equals $R_{B}(I;p_{M},\mathcal{Y}^{\prime}_{I})$ assuming $r(y,y;I)\equiv 0$).The Bayesian perspective provides additional insights to the works of Johnson et al., (2024); Ahdritz et al., (2024).
+It is clear that (5) matches these behaviours at the extremes: it equals 0 in the former case, and in the latter case correctly indicate a fully reducible risk.
+The Bayesian perspective provides additional insights to the works of Johnson et al., (2024); Ahdritz et al., (2024).
 Eq.
 (5) generalizes their methodology by allowing for a wider range of choices for $\{x_{n+1:N}\}$, and the above discussion provides guidance on their choices.
-Finally, we note that while (7) is generally applicable to MBR decoding, with a trivial choice of $\mathcal{Y}_{I}^{\prime}$ it may take a trivial value: if we use the Gibbs predictor which always samples from $p_{M}(\cdot\mid x_{1:n},y_{1:n})$, we will find (7) to be zero; this is to be expected as the predictor does not make any attempt to utilize the additional information.
-In our experiments we choose $\mathcal{Y}_{I}^{\prime}$ to contain multiple samples, one from each sample path for $y_{n+1:N}$.
-As we will see, this is sufficient to provide meaningful estimates of epistemic uncertainty.
+Remark 2.2 (generalizations).
+(5) can be readily applied beyond ICL if we use $(x_{n+1:N},y_{n+1:N})$ to represent more general queries, e.g., any clarifying questions and their answers.
+In such scenarios, (5) remains a non-negative measure of reducible uncertainty.
+While we will focus on ICL in experiments, we note that this generalization can be useful when the real responses are more expensive to obtain compared to LM generations, for example when they require interaction with user (Li et al.,, 2023) or a third-party source (Min et al.,, 2024).
+3 Related Work
+Probabilistic UQ for LMs.
+UQ has long been studied in LM applications involving a short or fixed-form response (e.g., Kamath et al.,, 2020; Jiang et al.,, 2021).
+Free-form generation introduces additional challenges.
+Our discussions are motivated by recent works on semantic uncertainty (Kuhn et al.,, 2023; Tian et al.,, 2023), which we seek to clarify and extend to a general scope.
+The measure (2) is a generalized entropy (DeGroot,, 1962); similar to Xu et al., (2019, p.8), we deviate from that line of classical work by considering general loss functions and computational constraints.
+All methods reviewed so far can be viewed as adopting a probabilistic perspective: they quantify uncertainty through the variation in the LM’s predictive distribution.
+Another line of work proposes to prompt the LM for its uncertainty, which we discuss shortly.
+Orthogonal to these developments, conformal prediction methods (e.g., Quach et al.,, 2023; Mohri and Hashimoto,, 2024; Yadkori et al., 2024b, ) take a prespecified uncertainty measure and a calibration set as input and provide a predictor with coverage guarantees.
+Calibration in free-form generation.
+Band et al., (2024) studied calibration in scenarios where the free-form generation $y$ is used for downstream classification based on a known predictor $\phi(y)$.
+Our work generalizes their setup, as we can always define a utility for $y$ based on $\phi$ and the standard utility functions for classification.
+Huang et al., 2024a studied the rank-calibration of general uncertainty measures which, as discussed in their work, is a different property.
+We note that an LM could be arbitrarily overconfident or underconfident for the uncertainty measure (2) to remain rank-calibrated.
+Huang et al., 2024b proposed several calibration metrics based on a similarity measure.
+In their work both the similarity measure and the method to aggregate calibration error are left underspecified, rendering the resulted metrics hard to interpret; in contrast, our Eq.
+(4) admits a straightforward interpretation as a generalized ECE.
+Prompt-based methods, instruction-tuned LMs.
+A recent line of work proposes to quantify uncertainty through prompting, e.g., by asking the LM to output the probability that its generation is correct (Xiao and Wang,, 2021; Mielke et al.,, 2022; Ren et al.,, 2023) or to answer a multiple-choice question (Kadavath et al.,, 2022).
+Compared with the probabilistic methods, this approach can be more efficient computationally as it avoids the need to draw multiple samples for each query.
+However, it requires the LM to have stronger reasoning capabilities, and is more difficult to apply when the user’s utility involves factors beyond factual correctness or for the quantification of epistemic uncertainty.
+Still, within the above scope, it is reasonable to ask whether the verbalized approach might be preferable for instruction-tuned LMs.
+The question remains largely open: previous studies have found that instruction tuning can affect probabilistic calibration (Achiam et al.,, 2023; Tian et al.,, 2023), but empirical comparisons between probabilistic and verbalized approaches yielded inconsistent results (Kuhn et al.,, 2023; Tian et al.,, 2023; Xiong et al.,, 2024).
+Moreover, in the free-form setting there is also a lack of principled evaluation of calibration: many prior works (Kuhn et al.,, 2023; Lin et al.,, 2024) adopted evaluation metrics that conflate calibration with predictive performance (Huang et al., 2024a, , p.5), and some studies (Jiang et al.,, 2021; Tian et al.,, 2023) evaluated token-level (rather than task-specific) calibration, which is less relevant.
+Our experiments in §4.1 will fill in this gap.
+Epistemic uncertainty.
+Our method is inspired by Fong et al., (2023); Xu and Raginsky, (2022).
+Similar ideas were discussed by Berti et al., (2021); Wen et al., (2022) in relation to Fong et al., (2023), and Grünwald and Dawid, (2004); Kotelevskii and Panov, (2024) in relation to Xu and Raginsky, (2022).
+Similar notions of reducible uncertainty also appear in active learning (Smith et al.,, 2023), and in experiment design (Rainforth et al.,, 2024) in relation to Remark 2.2.
+All of these works either restricted to the estimation of a full data distribution, or required explicitly Bayesian models,333 i.e., a $p_{M}$ that corresponds to a Bayesian posterior predictive, defined by a known posterior over model parameters and a likelihood.
+rendering them inapplicable to general LM applications.
+Falck et al., (2024) applied the method of Fong et al., (2023) to ICL, but similarly restricted to density estimation for numerical data and approximately Bayesian models.
+Concurrent work of Yadkori et al., 2024a proposed a measure similar to (5), but based on the log loss.
+It thus cannot distinguish between task-specific and irrelevant uncertainties, and also requires access to the LM’s predictive density; our work avoids such limitations.
 4 Experiments
-3.1 Free-Form Question Answering
-We first illustrate the proposed method on two QA tasks chosen by Lin et al., (2024): CoQA (Reddy et al.,, 2019) and NQOpen (Lee et al.,, 2019).
-CoQA contains reading comprehension questions where the answer can generally be deduced from the prompt, whereas NQOpen is an open-domain task and requires information not present in the prompt.
-Kalai and Vempala, (2024) proved in simplified scenarios that instruction-tuned LMs may have to be uncalibrated on the latter type of task; we investigate their theoretical predictions empirically.
-Figure 1: Reliability diagrams for GPT-3.5 on QA tasks. Error bar denotes 95% confidence interval for the average observed utility.
-We follow the setup in Lin et al., (2024) and adopt the GPT-3.5 model (OepnAI,, 2024), both for generation and to define the automated evaluation criterion; the latter determines our utility function and in turn the uncertainty measure (2).444 In this aspect our method differs from Lin et al., (2024), where uncertainty measures are defined with estimated entailment relations or lexical metrics.
-The evaluation criterion is defined using ICL, by instructing the LM to rate the consistency between two responses based on few-shot demonstrations (Lin et al.,, 2024, App.
-B.3).
-We reuse the model samples released by Lin et al., (2024), which contains 20 sampled responses for each query.
-Following their work we focus on the Gibbs predictor.
-The results are shown in Fig.
-1.
-As we can see, the LM’s subjective uncertainty generally correlates well with the average observed utility.
-It is slightly overconfident on the reading comprehension task, but much more so on the open-domain task.
-Such a difference appears consistent with the theoretical analysis of Kalai and Vempala, (2024).
-3.2 In-Context Machine Translation
-We now consider machine translation with ICL, where utility is defined through the widely adopted chrF score (Popović,, 2015).
-chrF measures both semantic and syntactic differences; the latter can be relevant in translation due to the need to detect unnatural or invalid syntactic structures.
-The inherent syntactic differences in the possible generations for $y\mid x$ constitute a source of irreducible uncertainty, making machine translation a non-trivial scenario to illustrate our uncertainty decomposition method.
-We follow the setup of Agarwal et al., (2024) who demonstrated that ICL can lead to steady improvements for machine translation.
-We adopt the FLORES+ dataset (NLLB Team et al.,, 2022), in particular the translation tasks from English to the following languages: French, Tamil, Yue Chinese (Cantonese), and Tigrinya.
-We use the following LMs: GPT-3.5, Gemini-1.5-Flash, and Gemini-1.5-Pro (Gemini Team,, 2024), and vary the number of ICL shots $n\in\{4,16\}$.
-For each task, we subsample 128 queries $(x_{*},y_{*})$ from the datasets.
-For each query, we draw 35 samples from the LM to approximate its predictive distribution, and another 5 samples to define the action space $\mathcal{Y}^{\prime}_{I}$.
-For the quantification of epistemic uncertainty, we set $N=n+8$, and sample $x_{n+1:N}$ by instructing an LM to rewrite the test query $x_{*}$; see Appendix B for details.
-$n$ gemini-1.5-flash gemini-1.5-pro gpt-3.5
-$4$ 0.276 / 0.356 0.332 / 0.431 0.185 / 0.329
-$16$ 0.246 / 0.371 0.304 / 0.426 0.168 / 0.351
-Table 1: Machine translation (English $\to$ Yue Chinese): gECE (Eq. (6), $\downarrow$) and chrF ($\uparrow$) for varying choices of $n$.
-We first evaluate the calibration of the combined subjective uncertainty on the Yue Chinese task.
-As shown in Table 1, the LMs are generally poorly calibrated, although calibration improves as the sample size $n$ increases.
-Calibration plots are presented in Figure 3 in Appendix, which shows that all LMs are overconfident.
-It is interesting to note an anticorrelation between the task performance and calibration among the three LMs evaluated.
-We note that Yue Chinese may constitute a low-resource translation task (Haddow et al.,, 2022) for which the LMs’ pretraining corpus is likely less informative.
-As is also shown in Figure 3, calibration is better on the French task for which more corpus is available.
-Figure 2: Machine translation (English $\to$ Yue Chinese): reduction of epistemic uncertainty (EU) vs total uncertainty (TU) as the ICL sample size increases from $4$ to $16$.
-We now investigate whether the LMs possess a coherent notion of epistemic and aleatoric uncertainty, in particular whether epistemic uncertainty is the only reducible component as sample size increases.
-For this purpose we plot the reduction of total uncertainty (TU, (2)) against that of epistemic uncertainty (EU, (7)) on each data point.
-Figure 2 shows that the two quantities have a good match.
-Deviation from perfect equivalence is expected here, due to the variance in the Monte-Carlo estimation for EU and TU.
-Language French Tamil Yue Tigrinya
-Avg. EU 0.016 0.059 0.078 0.091
-chrF 0.737 0.392 0.329 0.084
-# speakers (M) 312 87 87 10
-Table 2: Machine translation: average epistemic uncertainty on different languages, obtained from gpt-3.5 using $n=4$ ICL shots.
-We also provide the chrF score and the number of speakers (from Ethnologue,, 2024, in million) of each language as an indication of its linguistic resource available.
-Finally, we evaluate if the epistemic uncertainty measure (7) is meaningful, by comparing its average value across different translation tasks: we expect sensible notions of epistemic uncertainty to be higher on tasks for low-resource languages, as they are likely less represented in the pretraining corpus.
-Table 2 confirms this intuition: the LM demonstrates higher epistemic uncertainty on lower-resource languages.
+4.1 Uncertainty in Question Answering
+We first illustrate the uncertainty and calibration measures by revisiting the question answering (QA) experiments in recent works (Lin et al.,, 2024; Tian et al.,, 2023).
+As discussed in §3, our goal is to determine whether probabilistic UQ methods are suitable for modern instruction-tuned LMs, by providing a principled evaluation of task-specific calibration.
+Experiment setup.
+We combine the QA datasets studied by Lin et al., (2024); Tian et al., (2023): CoQA (Reddy et al.,, 2019), TriviaQA (Joshi et al.,, 2017), NQOpen (Kwiatkowski et al.,, 2019), SciQ (Welbl et al.,, 2017), and TruthfulQA (Lin et al.,, 2021).
+We subsample 1000 queries for each dataset and report bootstrap confidence intervals (CIs) for all metrics.
+The utility $S(y^{\prime},y;I)$ is defined by instructing an LM to rate the consistency of $y^{\prime}$ with $y$ as an answer to the user’s question.
+We adopt the prompt templates in Lin et al., (2024) for evaluation and generation.
+We add a system prompt that instruct LMs to generate single-line responses, and use gpt-4o-mini for evaluation.444 We confirmed that evaluation remains aligned with human perception, and that except on SciQ the generations are still varied in form.
+SciQ is a short-form dataset which we include for consistency with Tian et al., (2023).
+As TriviaQA and NQOpen contain trivia questions that were annotated well before the knowledge cutoff date of current LMs, we further remove a subset of questions with clearly outdated answers; see App.
+B.1 for details.
+We use LMs from the following model families: gpt-4o (OpenAI,, 2024), llama-3.1 (Dubey et al.,, 2024), gemini-1.5 (Gemini Team,, 2024), and claude-3.5 (Anthropic,, 2024).
+All LMs are instruction tuned unless otherwise noted.
+In a subset of experiments, we compare the uncertainty measure (2) (denoted as Prob.) with the prompt-based methods of Kadavath et al., (2022, denoted as P(True)) and Tian et al., (2023, Verb.).
+We apply all methods to the Gibbs predictor, and estimate uncertainty measures using Monte-Carlo estimation with 10 samples for each query.
+To evaluate calibration, we use the ECE metric (4) estimated through histogram binning (Naeini et al.,, 2015).
+For reference, we also compute the average test utility to quantify predictive performance, as well as the AURAC metric from Nadeem et al., (2009); Lin et al., (2024), which measures the performance of a selective prediction procedure based on a given LM and uncertainty measure.
+Full setup details are deferred to App.
+B.1.
+Results and discussion.
+We defer full results to App. B.2 and summarize the main findings below.
+Figure 1: Question answering: ECE and utility for the llama-3.1-70b model before and after instruction tuning. Error bar denotes 95% bootstrap CI.
+How does instruction tuning affect calibration?
+We revisit an experiment in Tian et al., (2023) on the impact of instruction tuning on token-level calibration.
+Fig.
+1 plots the ECE and test utility of llama-3.1-70b before and after instruction tuning.
+Its left subplot can be compared with Tian et al., (2023, Fig.
+1 left, for the first three datasets).
+The comparison reveals a rather different phenomenon: whereas the calibration of token-level probability deteriorated significantly (an increase in ECE by 0.10.10.10.10.20.20.20.2) across all three datasets in Tian et al., (2023), we find the deterioration of task-specific calibration to be much smaller, except on TruthfulQA.
+It is interesting to note that among these datasets, TruthfulQA is the only one where instruction-tuning led to a significant improvement in utility; additional results on NQOpen and CoQA further confirm this trend.
+These results indicate that the impact of instruction tuning on calibration may have been overestimated in previous work.
+Results for AUARC are provided in App.
+B.2, where we also evaluate additional LMs.
+We find that across all datasets, the performance of selective prediction based on task-specific uncertainty is either approximately unchanged or improved after instruction tuning.
+This is similarly in contrast to Tian et al., (2023, Fig.
+1 right), which shows that selective prediction based on token-level uncertainty consistently performs worse after instruction tuning.
+Figure 2: Question answering: test utility vs ECE for isntruction-tuned LMs, averaged over the datasets in Lin et al., (2024). App. B.2 presents results for individual datasets.
+We now compare the calibration of instruction-tuned LMs with their predictive performance.
+As shown in Fig.
+2, calibration is generally correlated with predictive performance, but substantial variation remains.
+It is thus reasonable to expect that the calibration of many instruction-tuned LMs may still be improved with different tuning methodologies.
+Figure 3: Question answering: calibration error from different methods. Error bar denotes 95% bootstrap CI.
+Comparing probabilistic and prompt-based uncertainty.
+We compare the calibration of the uncertainty measure (2) with prompt-based alternatives, for which we can define an analogous ECE metric.
+Fig.
+3 visualizes ECE from different methods.
+We can see that the measure (2) generally leads to the best calibration error; the only exception is on TruthfulQA, which is “adversarially constructed” to evaluate LMs’ tendency to repeat human falsehoods.
+Reliability diagrams in App.
+B.2 provide further insights.
+Taken together, the results suggest that the probabilistic approach to UQ remains effective for modern instruction-tuned LMs.
+Note that in the above comparison, only the probabilistic uncertainty measure makes use of the utility $S$.
+However, as discussed before, in many scenarios it is reasonable to assume access to (a surrogate of) $S$.
+Moreover, the ability of the probabilistic approach to incorporate different choices of utility can also be viewed as an advantage, especially when utility cannot be interpreted as a probability of correctness.
+Figure 4: Machine translation: average utility vs the number of deferrals to the many-shot predictor.
+We also report the p-value of a permutation test that compares the AUC-DF from the EU method to random.
+4.2 Epistemic Uncertainty in ICL
+We now turn to epistemic uncertainty (EU) quantification in ICL, and investigate whether the proposed EU measure may provide an effective deferral strategy that routes selected queries to a more expensive predictor.
+Background and setup.
+LM generations can often be improved given information retrieved from an external source.
+Such a process can become expensive if retrieval is from a third party (Min et al.,, 2024) or if it leads to a very long prompt (Agarwal et al.,, 2024).
+In such scenarios, it would be desirable to limit retrieval to queries with a higher level of reducible uncertainty.
+Agarwal et al., (2024) showed that many-shot ICL can be effective in low-resource machine translation (Haddow et al.,, 2022).
+Informed by their findings, we evaluate the use of EU for choosing between a LM predictor given few-shot demonstrations and a more expensive predictor based on many-shot ICL.
+We compare this method (denoted as EU) to a random baseline, and an alternative based on total uncertainty (TU, Eq.
+(2)).
+We follow the setup in Agarwal et al., (2024) and adopt the chrF score (Popović,, 2015) as the utility.
+chrF is a lexical similarity score that measure both semantic and syntactic differences; the latter provides a source of irreducible uncertainty.
+We define a base predictor using ICL with $n=4$ shots, and a more expensive predictor with $n^{\prime}=128$.
+We defer to the expensive predictor on $m$ test queries with the highest EU, and report the average utility (over all test samples) w.r.t.
+$m$.
+Following Gupta et al., (2024), we refer to this function as a deferral curve, and use its integral (AUC-DF) as a scalar summary.
+The EU measure (5) is instantiated with $N-n=4$ and estimated as follows: for a query $I=\langle z_{1:n},x_{*}\rangle$, we first draw $l$ i.i.d.
+samples $\{z^{(i)}_{n+1:N}=(x^{(i)}_{n+1:N},y^{(i)}_{n+1:N})\}_{i=1}^{l}$ where $x^{(i)}_{n+1:N}\sim p_{x}(\cdot|I)$ and $y^{(i)}_{n+1:N}\sim\prod_{k=n+1}^{N}p_{M}(y_{k}\mid\langle z_{1:n},z_{n+1:k-1}^{(i)},x_{k}^{(i)}\rangle)$.
+Then, for each $i$, we sample $\{y^{(i,j)}_{*}\}_{j=1}^{m}\sim p_{M}(y_{*}\mid\langle z_{1:n},z_{n+1:N}^{(i)},x_{*}\rangle)$.
+We use $l(m-1)$ samples to define a plug-in Monte-Carlo estimator, and the remaining samples, $\{y^{(i,m)}_{*}\}_{i=1}^{l}$, to define the action space in (5).
+We define $p_{x}$ by using an LM to rewrite $x_{*}$, and choose $l=5,m=8$.
+Following Agarwal et al., (2024), we adopt the translation tasks from English to Kurdish (kmr) and Tamil (tam) in the FLORES+ dataset (NLLB Team et al.,, 2022).
+Due to resource limitations, we subsample 500 data points from each dataset.
+We consider the following LMs: gpt-4o-mini, gpt-4o and gemini-1.5-pro.
+Full setup details are deferred to Appendix B.3.
+Result and discussion.
+Fig.
+4 plots the deferral curves from the GPT models.
+We can see that the EU method consistently outperforms the random baseline, and the improvement in AUC-DF is statistically significant.
+In contrast, the TU baseline is not always better than random, which is expected since a larger number of in-context demonstrations will not help if predictive uncertainty is irreducible.
+To further compare between the EU and TU methods, we compute the bootstrap distributions of their AUC-DF metrics, which are obtained by resampling the test queries without replacement.
+As shown in Table 1, for all but one experiment (gemini-1.5-pro on kmr) EU outperforms TU in more than 75% of bootstrap simulations.
+Table 1: Machine translation: AUC-DF difference between EU and TU methods. We report the median, $25\%$ and $75\%$ percentiles across 1000 bootstrap simulations.
+gemini-1.5-pro gpt-4o-mini gpt-4o
+tam $0.05$ $[0.00,0.09]$ $0.09$ $[0.01,0.16]$ $0.12$ $[0.06,0.19]$
+kmr $0.02$ $[-0.02,0.06]$ $0.15$ $[0.10,0.21]$ $0.15$ $[0.08,0.22]$
+In the gemini-1.5-pro experiments, both EU and TU demonstrate less improvement over the random baseline.
+This may be attributable to the fact that uncertainty from gemini-1.5-pro is less calibrated: the respective base predictor achieves an average ECE of 0.303 on the two datasets, compared with 0.104 using gpt-4o-mini and 0.083 using gpt-4o.
+Nonetheless, EU still appears to be the more effective method, as indicated by Table 1.
+See App.
+B.4 for full results.
 5 Conclusion
 This work studies uncertainty quantification in free-form natural language generation.
-Through a Bayesian decision-theoretic perspective, we derived principled methods to quantify the model’s task-specific subjective uncertainty and to evaluate its calibration.
-The proposed methods are also connected to previous work, which provides additional justification for their adoption.
-While the discussions (in particular those in §2.2 and §2.3) are not necessarily novel outside the context of language modelling, within this domain the decision-theoretic perspective appears new and addresses important conceptual challenges in uncertainty quantification.
-Experiments demonstrate that the methods extract meaningful uncertainty estimates from black-box LMs and quantify their calibration.
-Our discussion concerns the quantification of subjective uncertainty.
-In applications where i.i.d.
-calibration samples are available, it would be beneficial to recalibrate (Marx et al.,, 2022) the uncertainty measures, which can be achieved through (6), or to incorporate the uncertainty measures in conformal prediction (Quach et al.,, 2023).
-As the capabilities of LMs continue to evolve, it is also conceivable that future-generation models may find use in “one-of-its-kind” applications where post-calibration and conformal approaches are inapplicable.
-In such scenarios, it would be crucial to have principled measures for subjective uncertainty, as well as the tools to dissect and understand the sources of this uncertainty, e.g., through the proposed decomposition into epistemic and aleatoric uncertainty.
-It would also be interesting for future work to investigate the consistency between the LM’s verbalized uncertainty (Kadavath et al.,, 2022) and the subjective uncertainty defined by its predictive distributions.
-Appendix A Proof for the Equivalence between (7) and (7)
+Through a decision-theoretic perspective we derived principled methods to quantify the model’s task-specific subjective uncertainty and to evaluate its calibration.
+These methods are also connected to previous work, which provides additional justification for their adoption.
+While the discussions are not necessarily novel outside the context of language modelling, within this domain the decision-theoretic perspective appears new and addresses important conceptual challenges in uncertainty quantification.
+Experiments demonstrated the practical utility of the proposed methods.
+Appendix A Deferred Proof
 Claim A.1.
@@
 ∎
-Appendix B Additional Experimental Details
-In both experiments we use the following versions of LM APIs: gpt-3.5-turbo-0125, gemini-1.5-flash-001, gemini-1.5-pro-001.
-The machine translation experiment is implemented as follows: for each test query $x_{*}$, we sample the ICL demonstrations $(x_{1:n},y_{1:n})$, a single set of $x_{n+1:N}$, and $5$ sets of $y_{n+1:N}\sim p_{M}(x_{1:N},y_{1:n})$; then for each path sample $y_{n+1:N}$ we draw $8$ samples from $p_{M}(y_{*}\mid x_{1:N},y_{1:N},x_{*})$.
-The action space is defined using $5\times 1$ samples from each path sample, and the expectations in (2) and (7) using the remaining $5\times 7$ samples.
-For each $x_{*}$, we reuse the same set of samples for $(x_{n+1:N},x_{1:n},y_{1:n})\mid x_{*}$ across all experiments.
-The completion inputs $x_{n+1:N}$ are generated using Gemini-1.5-Flash with the following prompt:
+Appendix B Experimental Details
+Code and model samples for all experiments are available at https://github.com/meta-inf/suq-nlg.
+B.1 Question Answering: Deferred Setup Details
+We adopt the prompts in Lin et al., (2024) for generation and evaluation; see their Appendix B.
+For evaluation, their prompt instructs the LM to “rate the level of consistency” between a provided answer and a reference based on in-context demonstrations.
+For generation, we add a system prompt that instructs the LM to generate single-line responses to avoid interference with the evaluation prompt template in Lin et al., (2024); we find that without the system prompt, the more recent LMs we have evaluated have a greater tendency to generate multi-line responses.
+The generations can still contain multiple sentence, and we find them to be varied in format and content.
+The NQOpen and TriviaQA datasets contain a number of outdated questions, for example about the current holder of a public office or the latest season of an ongoing TV series.
+We thus use an LM (gpt-4o) to identify and remove such questions.
+We provide the LM with a question and its (possibly outdated) reference answer, and asks it to evaluate whether the question may fall into the above categories; the LM is asked to deprioritize its internal knowledge over the reference answer during evaluation.
+The prompt is shown in Fig.
+5.
+We note that the filtering process does not affect any finding reported in the main text; see e.g.
+Fig.
+6 (a) and (b).
+we chose to apply this process so that the reported results more accurately reflect the LMs’ performance on these tasks.
+We access all LMs through API services, from OpenAI (for GPT models) or Google Cloud.
+The API versions are as follows: gpt-4o-2024-08-06, gpt-4o-mini-2024-07-18, gemini-1.5-flash-001, gemini-1.5-pro-001, claude-3.5-sonnet@20240620 (versioning from Google cloud).
+We use the default sampling hyperparameters which always include a temperature of 1.
+We implement the prompt-based uncertainty measures as follows: for the method of Kadavath et al., (2022) we adopt their prompt template and include model generations in the prompt, as recommended in their work; for the method of Tian et al., (2023), we use the two-stage variant (Verb.
+2S top-1) which separates answer and confidence generation, so that the answer generation process can be comparable with the other methods; we use their prompt for confidence generation.
+For all test metrics we report bootstrap CIs computed using 10000 simulations.
+All prompts and an illustrative subset of model generations can be found at the code repository.
+I will provide you with a trivia question and its reference answer as of {year}. Please evaluate if the answer to this question is likely to have changed between {year} - 2023.
+Examples of questions that might have changed answers include:
+* Recent events or records
+* The latest season of an ongoing TV series
+* Current holders of public offices or titles
+Use your best judgment in the evaluation.
+In scenarios where your internal knowledge could be inaccurate, rely on common sense and the provided reference answer.
+Please format your response in two lines, as follows:
+1. A brief explanation of your reasoning (one sentence)
+2. A "Yes" or "No" answer indicating whether the answer likely changed
+Please adhere to the format of the example below:
+Question: when did the botswana currency first come into circulation
+Reference Answer: 1976
+Reasoning: The date a currency first entered circulation is a historical fact that doesnt change.
+Answer: No
+Now please answer the following:
+Question: {question}
+Reference Answer: {answer}
+Figure 5: Question answering: prompt template used to filter trivia datasets.
+B.2 Question Answering: Additional Results
+Full results in the setting of Fig.
+13 are provided in Fig.
+68, respectively, where we also provide results for AUARC.
+As expected, across LMs with a similar level of predictive performance, AUARC is generally correlated with calibration error: see the comparison between llama-3.1-70b and gemini-1.5-pro (or gpt-4o-mini) in Fig.
+7.
+It is also interesting to note that in the setting of Fig.
+1, the calibration of the gemma2-27b model (Gemma Team,, 2024) often improves after instruction tuning; inspection on TriviaQA and SciQ appears to show that the instruction-tuned LM has a similar level of variability in its generations, but the quality of generation has generally improved.
+Fig.
+9 compares the reliability diagrams obtained from different uncertainty measures, evaluated on llama-3.1-70b and gpt-4o.
+We can see that all uncertainty measures can be overconfident on inputs with low uncertainty, while the prompt-based measures can also be underconfident on high-uncertainty inputs.
+(a) llama-3.1-8b
+(b) gemma2-27b
+(c) llama-3.1-70b
+(d) llama-3.1-70b, without filtering for trivia datasets
+Figure 6: Question answering: full results in the setting of Fig. 1.
+Figure 7: Question answering: full results in the setting of Fig. 2. Error bar denotes 95% bootstrap CI.
+Figure 8: Question answering: results for AUARC in the setting of Fig. 3.
+(a) llama-3.1-70b, NQOpen
+(b) gpt-4o, NQOpen
+(c) llama-3.1-70b, TriviaQA
+(d) gpt-4o, TriviaQA
+(e) llama-3.1-70b, CoQA
+(f) gpt-4o, CoQA
+(g) llama-3.1-70b, SciQ
+(h) gpt-4o, SciQ
+(i) llama-3.1-70b, TruthfulQA
+(j) gpt-4o, TruthfulQA
+Figure 9: Question answering: reliability diagrams for the gpt-4o and llama-3.1-70b models. Note that the $x$-axis is always equivalent to the negated uncertainty measure.
+B.3 Machine Translation: Setup Details
+We use the prompt in Agarwal et al., (2024) for the translation task, and generate $x_{n+1:N}$ using gemini-1.5-flash with the prompt in Fig. 10. Details for the LMs involved are the same as App. B.1.
+The deferral experiments involve two ICL predictors.
+The base predictor is implemented by sampling $n=4$ demonstrations uniformly without replacement.
+For the more expensive predictor we append $n^{\prime}-n=124$ demonstrations that are closest to the test query in an embedding space, where embeddings are computed using the text-embedding-004 model from Google Cloud.
+[^6] https://cloud.google.com/vertex-ai/generative-ai/docs/model-reference/text-embeddings-api In both cases we retrieve demonstrations from the dev split of the dataset and evaluate on the devtest split.
+We experimented with both (nontrivial) MBR and Gibbs decoding; [^7] Note that these predictors do not coincide with the two predictors in (5).
+Instead, (5) is only used as a surrogate to approximate the performance improvement in the deferral process.
+the former is implemented by drawing $16$ samples for each query, using half of them to define $\mathcal{Y}^{\prime}_{I}$ and half to approximate the expectation in (1).
+The main text reports the results for the MBR predictors, as we find them to perform better across all experiments.
+However, the same EU measure also works well for the Gibbs predictors, as shown in Fig.
+12.
+Following Robinson et al., (2023), we use the chrF2++ score implemented by Post, (2018). The p-values reported in Fig. 4 are calculated using 10000 replications for the random baseline.
 Please generate {K} sentences that convey a broadly similar meaning as the one provided, ensuring they are of similar length and use varied wording.
@@
 {x_*}
-In the above, $x_{*}$ denotes the test query.
-For simplicity, we report the chrF scores evaluated using the final samples $y_{*}\mid x_{1:N},y_{1:N},x_{*}$, which is similar to the reinforced ICL method (Agarwal et al.,, 2024).
-Figure 3: Machine translation: reliability diagrams for a subset of experiments. Error bar indicates 95% confidence interval for the expected observed utility.
+Figure 10: Machine translation: prompt used to define $p_{x}$. $x_{*}$ denotes the test query.
+B.4 Machine Translation: Additional Results
+Table 2 reports the calibration error from all models. We can see that gemini-1.5-pro consistently demonstrates worse calibration.
+Table 2: Machine translation: ECE for all models.
+gemini-1.5-pro gpt-4o-mini gpt-4o
+Tam 0.250 0.041 0.065
+Kmr 0.356 0.017 0.010
+(a) English $\to$ Tamil
+(b) English $\to$ Kurdish
+Figure 11: Machine translation: results for gemini-1.5-pro in the setting of Fig. 4.
+(a) gemini-1.5-pro
+(b) gpt-4o-mini
+(c) gpt-4o
+Figure 12: Machine translation: deferral curve for Gibbs predictors on the English $\to$ Kurdish task.